### v1

# Armijo Line-search Makes (Stochastic) Gradient Descent Go Fast

Abstract
Armijo line-search (Armijo-LS) is a standard method to set the step-size for gradient descent (GD). For smooth functions, Armijo-LS alleviates the need to know the global smoothness constant $L$ and adapts to the “local” smoothness, enabling GD to converge faster. However, existing theoretical analyses of GD with Armijo-LS (GD-LS) do not characterize this fast convergence. We show that if the objective function satisfies a certain non-uniform smoothness condition, GD-LS converges provably faster than GD with a constant $1/L$ step-size (denoted as GD(1/L)). Our results imply that for convex losses corresponding to logistic regression and multi-class classification, GD-LS can converge to the optimum at a linear rate, and hence, improve over the sublinear convergence of GD(1/L). Furthermore, for non-convex losses satisfying gradient domination (for example, those corresponding to the softmax policy gradient in RL or generalized linear models with a logistic link function), GD-LS can match the fast convergence of algorithms tailored for these specific settings. Finally, we prove that under the interpolation assumption, for convex losses, stochastic GD with a stochastic line-search can match the fast convergence of GD-LS.
Gradient Descent, Armijo Line-search, Non-uniform Smoothness, Interpolation
1 Introduction
Gradient descent (GD) (Cauchy et al., 1847) and its stochastic variants are the preferred optimization methods in machine learning. The practical effectiveness of gradient methods heavily relies on the choice of the step-size (“learning rate”) parameter. Backtracking Armijo line-search (Armijo, 1966; Nocedal & Wright, 2006) (referred to as Armijo-LS) is a standard method to set the step-size for gradient descent. Given an initial step-size, the simplest form of Armijo-LS “searches” for the largest step-size that guarantees a sufficient decrease in the function value. When minimizing $L$-smooth functions using GD, Armijo-LS alleviates the need to know $L$, the global smoothness constant and enables setting the GD step-size in an adaptive manner. For both $L$-smooth convex and non-convex functions, GD with Armijo-LS (henceforth GD-LS) has been shown to match the favorable theoretical guarantees of GD with a constant $\nicefrac{{1}}{{L}}$ step-size (henceforth GD(1/L)). However, empirically, GD-LS typically results in faster convergence and is consequently, the default choice in practice.
One often-cited reason to explain the faster convergence of GD-LS is that it adapts to the “local” smoothness constant $L(\theta)$ near the point $\theta$, and results in an effective step-size of $\nicefrac{{1}}{{L(\theta)}}$. In some regions, $L(\theta)$ might be much smaller than the global smoothness $L$, thus allowing GD-LS to use much bigger step-sizes and consequently lead to faster convergence. However, existing theoretical analyses of GD-LS do not formalize this intuition, and can therefore, not explain the algorithm’s faster convergence.
In this paper, we aim to solve $\min_{\theta\in\mathbb{R}^{d}}f(\theta)$ for a special class of objective functions and formally characterize the advantage of GD-LS over GD(1/L). In particular, we make the following contributions.
Contribution 1. In Section 2, we introduce a class of functions that satisfy an $(L_{0},L_{1})$ non-uniform smoothness condition. This condition is similar to that proposed in Zhang et al. (2019) to explain the success of normalization and gradient clipping techniques when training neural networks. In particular, we consider functions where the local smoothness constant around a point $\theta$ is given by $L(\theta)=L_{0}+L_{1}\,f(\theta)$ where $L_{0}$ and $L_{1}$ are non-negative constants ($L_{1}=0$ corresponds to the standard uniform smoothness). We show that the proposed condition is satisfied by common objectives; for example, both the logistic and exponential losses used for linear classification satisfy the condition with $L_{0}=0$ and $L_{1}\neq 0$. Furthermore, we prove that this condition is also satisfied by non-convex functions corresponding to generalized linear models with a logistic link function and the softmax policy gradient objective in reinforcement learning.
Contribution 2. In Section 3, we analyze the convergence of GD-LS on functions satisfying the proposed conditions. For this, we first prove that the step-size selected by Armijo-LS around $\theta$ is lower-bounded by $1/L(\theta)$, and hence, GD-LS provably adapts to the local smoothness. We use this property to prove Theorem 1, a meta-theorem that quantifies the convergence rate of GD-LS. We note that Hübler et al. (2024) also consider GD-LS for minimizing a different class of non-uniform smooth functions. However, their analysis for general non-convex functions does not demonstrate the algorithm’s adaptivity to the smoothness, nor does it result in a faster rate than GD(1/L). Moreover, their resulting algorithm requires the knowledge of the non-uniform smoothness constant, making it impractical.
Contribution 3. In Section 4, we instantiate Theorem 1 for non-uniform smooth, convex losses that include logistic regression and multi-class classification with the cross-entropy loss as examples. Specifically, in Corollary 1, we show that GD-LS converges at an $O\left(\left(\nicefrac{{f^{*}}}{{\epsilon}}\right)\,\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ rate where $f^{*}:=\inf f(\theta)$. Hence, when $f^{*}$ is $O(\epsilon)$, GD-LS converges at an $O\left(\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ rate, compared to the sublinear $O\left(\nicefrac{{1}}{{\epsilon}}\right)$ convergence of GD(1/L). We instantiate this result for logistic regression on linearly separable data, and prove the linear convergence of GD-LS (Corollary 7), thus matching the rate for normalized GD (Axiotis & Sviridenko, 2023).
Contribution 4. In Section 5, we instantiate Theorem 1 for non-convex functions satisfying non-uniform smoothness and gradient domination conditions that guarantee global optimality. Specifically, in Section 5.1, we analyze the convergence of GD-LS on the softmax policy gradient objective in reinforcement learning (Mei et al., 2020). In this setting, the linear convergence rate attained by GD-LS is provably better than the $\Omega(\nicefrac{{1}}{{\epsilon}})$ convergence of GD(1/L) and matches the rate of natural policy gradient (Kakade & Langford, 2002). In Section 5.2, we analyze the convergence of GD-LS for functions satisfying the PL condition (Polyak, 1987; Karimi et al., 2016) and instantiate the result for generalized linear models with the logistic link function. Our result demonstrates that GD-LS can converge faster than both GD and normalized GD (Hazan et al., 2015).
Contribution 5. Finally, in Section 6, we show that the advantages of line-search carry over to the stochastic setting. Specifically, we consider a finite-sum objective $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}\,f_{i}(\theta)$, and study the convergence of stochastic gradient descent (SGD) in conjunction with a stochastic line-search (Vaswani et al., 2019b). We restrict our attention to the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt & Roux, 2013) which implies that each $f_{i}$ is minimized at $\theta^{*}:=\operatorname*{arg\,min}f(\theta)$. Such a condition is satisfied by logistic regression on linearly separable data. Interpolation enables the fast convergence of SGD, allowing it to match the GD rate but with an $O(1)$ iteration cost. Under interpolation, SGD with a stochastic line-search (referred to as SGD-SLS) and its variants empirically outperform constant step-size SGD, and have been used to train deep neural networks (Vaswani et al., 2019b; Galli et al., 2024). We provide further theoretical justification for the fast convergence of SGD-SLS. Specifically, in  Corollary 5, we prove that for logistic regression on linearly separable data, SGD-SLS converges at a linear rate.
2 Problem Formulation
We aim to solve the unconstrained minimization problem: $\min_{\theta\in\mathbb{R}^{d}}f(\theta)$. We define $\theta^{*}\in\arg\inf f(\theta)$ as an optimal solution and $f^{*}:=\inf f(\theta)$ as the minimum function value. Throughout, we consider $f$ to be twice-differentiable and satisfies the following assumptions:
Assumption 1.
$f$ is non-negative i.e. for all $\theta$, $f(\theta)\geq 0$.
Assumption 2.
$f$ is $(L_{0},L_{1})$ non-uniform smooth i.e. for $L_{0},L_{1}\geq 0$,
(a) For all $x,y$ such that $\left\|x-y\right\|\leq\frac{q}{L_{1}}$ where $q\geq 1$ is a constant, if $A:=1+e^{q}-\frac{e^{q}-1}{q}$ and $B:=\frac{e^{q}-1}{q}$,
$f(y)$ $\leq f(x)+\langle\nabla f(x),y-x\rangle$
$+\frac{(A\,L_{0}+B\,L_{1}\,f(x))\,}{2}\left\|y-x\right\|_{2}^{2}$ (1)
(b) For all $\theta$, $\left\|\nabla^{2}f(\theta)\right\|\leq L_{0}+L_{1}\,f(\theta)$
If $L_{1}=0$Assumption 2 recovers the standard uniform smoothness condition as a special case. Consequently, common smooth objectives such as linear regression or logistic regression satisfy the above condition. For example, if $X\in\mathbb{R}^{d\times n}$ is the feature matrix, and $y\in\mathbb{R}^{n}$ is the vector of measurements, then the linear regression objective $f(\theta)=\frac{1}{2n}\,\left\|X\,\theta-y\right\|_{2}^{2}$ is $(\frac{1}{n}\lambda_{\max}[X^{T}X],0)$ non-uniform smooth where $\lambda_{\max}[A]$ is the maximum eigenvalue of the PSD matrix $A$. The above condition is similar to the non-uniform smoothness conditions proposed in the literature (Zhang et al., 2019, 2020; Chen et al., 2023). The difference is that the smoothness in Assumption 2 is proportional to $f(\theta)$ rather than $\left\|\nabla f(\theta)\right\|$ in the previous work. Subsequently, we will see that GD with Armijo line-search is easier to analyze with this alternate definition of non-uniform smoothness.
In order to show the benefit of GD with Armijo line-search, we focus on functions where $L_{1}\neq 0$. We will require these functions to satisfy an additional assumption that relates the gradient norm to the function value. As we will see, such an assumption is typically true for losses with an exponential tail, even when using a 2 layer neural network (Taheri & Thrampoulidis, 2023; Wu et al., 2024).
Assumption 3.
For all $\theta$, there exist constants $\omega,\nu>0$ s.t.
$\left\|\nabla f(\theta)\right\|\leq\nu\,f(\theta)+\omega.$ (2)
To motivate the above assumptions, we prove that common convex objectives for supervised learning such as linear logistic regression and linear multi-class classification satisfy the above condition with $L_{0}=0$ and non-zero $L_{1}$. Moreover, we show that these functions also satisfy Assumption 3 (all proofs are deferred to Appendix A). Below, we state the result for logistic regression and defer the results for the other losses to the Appendix.
Proposition 1.
Consider $n$ points where $x_{i}\in\mathbb{R}^{d}$ are the features and $y_{i}\in\{-1,1\}$ are the corresponding labels. Logistic regression with the objective
$f(\theta)=\frac{1}{n}\,\sum_{i=1}^{n}\ln(1+\exp(-y_{i}\langle x_{i},\theta\rangle)$ (3)
satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$, and Assumption 3 with $\nu=\max_{i}\left\|x_{i}\right\|$ and $\omega=0$.
Note that the logistic regression objective is also uniform smooth, meaning that it simultaneously satisfies Assumption 2 with $L_{0}=\frac{1}{4n}\,\lambda_{\max}[X^{T}X]$ and $L_{1}=0$, where $X\in\mathbb{R}^{n\times d}$ is the corresponding feature matrix. On the other hand, binary classification with the exponential loss is not uniform smooth on an unbounded domain, but satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ (see Proposition 5 in Appendix A). We note that for logistic regression, the loss corresponding to a single point $i\in[n]$ satisfies the notion of non-uniform smoothness in Zhang et al. (2019) with $L_{0}=0$ and $L_{1}=\left\|x_{i}\right\|$ (Gorbunov et al., 2024, Example 1.6). However, $f(\theta)$ does not necessarily satisfy this assumption with $L_{0}=0$ (see Proposition 7 for a simple counter-example). Subsequently, we will see that having $L_{0}=0$ is key to achieving fast convergence for GD-LS. This further motivates our alternate definition of non-uniform smoothness.
Next, we show that the non-convex objective corresponding to generalized linear models (GLM) with a logistic link function also satisfies Assumptions 1, 2 and 3 . In particular, we prove the following proposition in Appendix A.
Proposition 2.
Consider $n$ points where $x_{i}\in\mathbb{R}^{d}$ are the features and $y_{i}\in[0,1]$ are the corresponding labels. If $\pi_{i}(\theta)=\sigma(\langle x_{i},\theta\rangle):=\frac{1}{1+\exp(-\langle x_{i},\theta\rangle)}$, the GLM objective
$f(\theta)=\frac{1}{2n}\sum_{i=1}^{n}\left(\pi_{i}(\theta)-y_{i}\right)^{2}$ (4)
satisfies Assumption 2 with $L_{0}=\frac{17}{16}\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and $L_{1}=2\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and Assumption 3 with $\nu=2\max_{i}\left\|x_{i}\right\|$ and $\omega=\max_{i}\left\|x_{i}\right\|$.
Finally, in Appendix A, we also show that the objective for softmax policy gradient (Mei et al., 2020) also satisfies the required assumptions for multi-armed bandits and tabular Markov decision processes, and study these settings in more detail in Section 5.
Now that we have motivated the use of consider Assumptions 1 and 3, in the next section, we will consider minimizing non-uniform smooth functions using gradient descent.
3 GD with Armijo Line-search
We consider using gradient descent (GD) with Armijo line-search  (Armijo, 1966) (henceforth referred to as GD-LS) to minimize functions satisfying Assumptions 1, 2 and 3. The GD-LS update at iteration $t\in[T]$ is given as:
$\theta_{t+1}$ $=\theta_{t}-{\eta_{t}}\,\nabla f(\theta_{t})\,,$ (5)
where ${\eta_{t}}$ is the step-size returned by backtracking Armijo line-search (Armijo-LS). In particular, starting from an initial maximum step-size $\eta_{\max}$, Armijo-LS uses backtracking to select a step-size that satisfies the Armijo condition,
$f(\theta_{t}-\eta_{t}\nabla f(\theta_{t}))$ $\leq f(\theta_{t})-c\eta_{t}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\,,$ (6)
where $c\in(0,1)$ is a tunable parameter. The complete pseudo-code is described in Algorithm 1. The parameter $\beta$ controls the backtracking and is typically set to $0.9$, while the parameter $c$ is typically set to a small value such as $10^{-4}$ (Nocedal & Wright, 2006).
1:  Input: $\theta_{0}$, $\eta_{\max}$, $c\in(0,1)$, $\beta\in(0,1)$
2:  for $t=0,\dots,T-1$ do
3:     $\tilde{{\eta_{t}}}\leftarrow$ $\eta_{\max}$
4:     while $f(\theta_{t}-\tilde{{\eta_{t}}}\,\nabla f(\theta_{t}))>f(\theta_{t})-c\,\tilde{{\eta_{t}}}\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ do
5:        $\tilde{{\eta_{t}}}\leftarrow\tilde{{\eta_{t}}}\beta$
6:     end while
7:     ${\eta_{t}}\leftarrow\tilde{{\eta_{t}}}$
8:     $\theta_{t+1}=\theta_{t}-{\eta_{t}}\nabla f(\theta_{t})$
9:  end for
Algorithm 1 GD with Armijo Line-search (GD-LS)
When using GD-LS for minimizing $L$ (uniformly)-smooth functions (corresponding to $L_{0}\neq 0$ and $L_{1}=0$ in Assumption 2), ${\eta_{t}}$ is constrained to lie in the $\left[\min\left\{\eta_{\max},\frac{2\,(1-c)\,\beta}{L}\right\},\eta_{\max}\right]$ range. Note that this bound holds for all $L$ (uniformly)-smooth functions, does not require convexity, and guarantees that backtracking line-search will terminate at a non-zero step-size. The parameter $c$ controls the “aggressiveness” of the algorithm; small $c$ values encourage a larger step-size. Hence, Armijo-LS can be seen as method to obtain a step-size proportional to $1/L$ without the knowledge of the global smoothness constant.
The bounds on the step-size can be used to derive the convergence rate for GD-LS. For example, for uniformly $L$-smooth and convex functions, the standard analysis shows that GD-LS converges to the optimum at an $O(1/T)$ rate (Nocedal & Wright, 2006), thus matching the rate of GD with a constant step-size equal to $1/L$ (henceforth referred to as GD(1/L)). However, as alluded to in Section 1, Armijo-LS enables GD to adapt to the “local” smoothness $L(\theta_{t})$ (the smoothness around iterate $\theta_{t}$), and results in faster convergence in practice. By studying non-uniform smooth functions satisfying Assumption 2, we aim to formally characterize this fast convergence.
To that end, we first show that when minimizing non-uniform smooth functions satisfying Assumption 2, Armijo-LS can result in provably larger step-sizes that enable faster convergence. For the subsequent theoretical analysis, we only consider “exact backtracking” i.e. we assume that the backtracking procedure returns the largest step-size that satisfies the Armijo condition, meaning that $\beta\approx 1$. It is straightforward to relax this assumption similar to the standard analysis (Nocedal & Wright, 2006). We first prove the following lemma on the minimum step-size returned by Armijo-LS.
Lemma 1.
If $f$ satisfies Assumptions 1, 2 and 3, at iteration $t$, GD-LS returns a step-size
${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}\right\}\,,$
where $\lambda_{0}:=3\,\frac{L_{0}+L_{1}\,\omega}{(1-c)}$ and $\lambda_{1}:=3\,\frac{L_{1}(\nu+1)}{(1-c)}$.
Proof Sketch.
For “smaller” step-sizes $\eta\leq\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$, we use Eq. 1 to prove that the step-size will satisfy the Armijo condition, and the backtracking line-search will terminate. However, since Eq. 1 is only valid when $\left\|y-x\right\|$ is small, we cannot directly use it when the step-size returned by Armijo-LS is “large”. For this, we first prove that $g(\theta):=\ln(f(\theta))$ is $L_{1}$ uniformly-smooth. We then show that the step-size returned by the line-search on $f(\theta)$ satisfies an equivalent Armijo condition on $g(\theta)$, and use the uniform smoothness of $g(\theta)$ to lower-bound the step-size. ∎
Note that when $L_{1}=0$, the lower-bound on ${\eta_{t}}$ is similar to that for uniformly smooth functions. However, when $L_{0}=0$ (e.g. for logistic regression or multi-class classification), the lower-bound on ${\eta_{t}}$ is proportional to $1/f(\theta_{t})$, meaning that as $f(\theta_{t})$ decreases, the step-size returned by Armijo-LS increases. This enables the faster convergence of GD-LS.
We now present a meta-theorem (proved in Appendix B) which we will instantiate and interpret for convex losses in Section 4 and non-convex losses in Section 5.
Theorem 1.
For a fixed $\epsilon>0$, if $f$ satisfies Assumptions 1, 2 and 3 and if for a constant $R>0$, $\left\|\nabla f(\theta)\right\|_{2}^{2}\geq\frac{[f(\theta)-f^{*}]^{2}}{R}$, ${f^{*}}>0$, then GD-LS with $\eta_{\max}=\infty$ requires
$T\geq\begin{cases}&\max\{2\,R\lambda_{1},1\}\,\left(\frac{{f^{*}}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-{f^{*}}}{\epsilon}\right)\\ &\text{ if $f^{*}\geq\frac{\lambda_{0}}{\lambda_{1}}-\epsilon$}\;\;\;\textbf{(Case (1))}\\[8.61108pt] &\frac{2\lambda_{0}\,R}{\epsilon}+\max\{2\,R\lambda_{1},1\}\,\left(\frac{{f^{*}}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-{f^{*}}}{\epsilon}\right)\\ &\text{ otherwise }\;\;\;\textbf{(Case (2))}\end{cases}$
iterations to ensure to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
Proof Sketch.
Using the condition $\left\|\nabla f(\theta)\right\|_{2}^{2}\geq\frac{[f(\theta)-f^{*}]^{2}}{R}$ with the Armijo condition in Eq. 6 and the lower-bound on ${\eta_{t}}$ from Lemma 1, we get that
$f(\theta_{t+1})\leq f(\theta_{t})-\frac{[f(\theta_{t})-f^{*}]^{2}}{[\lambda_{0}+\lambda_{1}\,f(\theta_{t})]\,R}$ (7)
We split the proof into two cases: Case (1) when $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$ for all $t\in[T]$ iterations required to obtain the desired sub-optimality. In this case, $\lambda_{1}\,f(\theta_{t})\geq\lambda_{0}$. Using this relation with Eq. 7 and following the arguments in Axiotis & Sviridenko (2023, Theorem 5.2) allows us to complete the proof of Case (1). For Case (2), we follow the proof of the first case for iterations $t\in[\tau]$ for which $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$, and obtain a similar rate. This corresponds to Phase 1. with faster convergence. For all iterations $t\in[\tau,T]$, $f(\theta_{t})\leq\frac{\lambda_{0}}{\lambda_{1}}$ and hence $\lambda_{0}\geq\lambda_{1}\,f(\theta_{t})$. Using this relation with Eq. 7 and following the standard proof for uniformly smooth functions completes the proof for Phase 2. which results in slower convergence. Putting together the results for both phases completes the proof of Case (2). ∎
We note that the above theorem requires an additional condition which lower-bounds the gradient norm in terms of the function sub-optimality. In Section 4, we use convexity to satisfy this condition, whereas in Section 5, we use gradient domination to satisfy it. We also require $\eta_{\max}$ to be larger than $\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$ for all $t$. This ensures that the step-size is not constrained by the initial choice, but rather by the properties of the function. For conciseness, we express this condition as $\eta_{\max}=\infty$, and note that it is straightforward to relax it, albeit at the cost of clarity.
In order to interpret the above theorem, let us first consider the setting corresponding to $\lambda_{1}=0$. Here, case (2) is active, and the algorithm requires $O(1/\epsilon)$ iterations to achieve the desired sub-optimality, matches the standard result for uniformly smooth functions. Now consider the setting when $\lambda_{0}=0$. Here, case (1) is active, and GD-LS requires $O\left(R\,\left(\frac{f^{*}}{\epsilon}\right)\,\ln\left(\frac{1}{\epsilon}\right)\right)$ iterations. The iteration complexity thus depends on $f^{*}$, and in cases where $f^{*}$ is small, GD-LS can result in an improved rate. As a concrete example, consider the case when $f^{*}=\delta\,\epsilon$ where $\delta\geq 1$ is a constant independent of $\epsilon$. In this setting, GD-LS will result in a faster $O\left(R\,\ln\left(\frac{1}{\epsilon}\right)\right)$ convergence. Note that GD(1/L) does not benefit from such adaptivity, and will always result in a sublinear rate. For non-zero $\lambda_{0}$ and $\lambda_{1}$, the resulting rate depends on the value of $f^{*}$. If $f^{*}$ is larger than the threshold $\frac{2\lambda_{0}}{\lambda_{1}}$, GD-LS can result in the potentially fast rate corresponding to case (1) whereas if $f^{*}$ is smaller than the threshold, the algorithm has a two-phase behaviour: fast convergence until the loss becomes smaller than the threshold followed by slow convergence to the minimizer.
In the next section, we instantiate the above theorem to prove the fast convergence of GD-LS for convex losses.
4 GD-LS for Convex Losses
In this section, we characterize the convergence rate of GD-LS on convex losses satisfying Assumptions 1, 2 and 3. Recall that binary classification using logistic regression and multi-class classification using the cross-entropy loss both satisfy Assumptions 1, 2 and 3 with $L_{0}=0$ and $\omega=0$. Consequently, we instantiate Theorem 1 for this setting and prove the following corollary in Appendix C.
Corollary 1.
For a fixed $\epsilon>0$, assuming $f(\theta)$ is convex and satisfies Assumptions 1, 2 and 3 with $L_{0}=0$ and $\omega=0$, GD-LS with $\eta_{\max}=\infty$, requires $T\geq$
$\max\{2\lambda_{1}\,\left\|\theta_{0}-\theta^{*}\right\|_{2}^{2},1\}\,\left(\frac{f^{*}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-f^{*}}{\epsilon}\right)$
iterations to ensure to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
Figure 1: Comparing GD-LS with $c=1/2$ and $\eta_{\max}=10^{8}$ and GD(1/L) for unregularized logistic regression on the ijcnn dataset (Chang & Lin, 2011). $f^{*}$ is small and GD-LS converges faster.
Referring to the explanation following Theorem 1, we conclude tht GD-LS can result in faster convergence over GD(1/L) when $f^{*}$ is small (see Fig. 1 for an experimental validation). In order to compare the result in Corollary 1 with existing works, let us consider the special case of logistic regression. In this case, GD-LS matches the rate for a variant of normalized gradient descent (NGD) in Axiotis & Sviridenko (2023, Theorem 5.2). However, unlike GD-LS, NGD requires the knowledge of $L_{1}$ making it relatively difficult and less likely to be implemented in practice. Furthermore, we note NGD is a specialized algorithm that is helpful to attain faster rates for certain problems (Mei et al., 2021; Hazan et al., 2015; Wilson et al., 2019), whereas GD-LS is universally used and can automatically exploit the problem structure. Moreover Corollary 1 is more general and also holds for the exponential loss.
While GD(1/L) can result in an $\Omega(1/\epsilon)$ rate for general smooth, convex functions (Nesterov et al., 2018), analyses of GD on logistic regression often exploit strong-convexity and prove faster rates (Karimi et al., 2016). In particular, if the iterates lie in a bounded set, the objective is $\mu(\theta)$ strongly-convex where $\mu(\theta)=\lambda_{\min}[X^{T}X]\,\min_{i}\pi_{i}(\theta)\,(1-\pi_{i}(\theta))$ where $\pi_{i}=\frac{1}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}$. Notice that as $\pi_{i}(\theta)$ tends to either zero or one i.e. the predictions become deterministic, $\mu(\theta)$ tends to 00Freund et al. (2018, Theorem 3.3) characterize the resulting rate for GD(1/L) and prove that the suboptimality scales as $O\left(\exp(-T\,\exp\left(\nicefrac{{-1}}{{\xi}}\right))\right)$ where $\xi$ is the degree of non-separability and tends to zero as the data becomes more separable. Hence, the rate becomes exponentially worse as $\xi$ decreases. However, as the data becomes separable, GD-LS converges at a faster linear rate (see Fig. 2), meaning that strong-convexity cannot explain this behaviour.
Consequently, we consider the special case when the data is linearly separable and $f^{*}=0$ and better characterize the fast convergence for logistic regression. Unfortunately, we cannot directly use Corollary 1 since $\left\|\theta\right\|\to\infty$ as $f(\theta)\to 0$, making the resulting bound vacuous (Orabona, 2024). Consequently, we use a different technique and first prove the following theorem in Appendix C.
Theorem 2.
For a fixed $\epsilon>0$ and an arbitrary comparator $u$, if $f(\theta)$ is convex, $L$-uniform smooth and satisfies Assumptions 1, 2 and 3 with $L_{0}=0,\,\omega=0$, GD-LS with $\eta_{\max}=\infty$ and $c>\frac{1}{2}$, requires
$T\geq\frac{c\,\lambda_{1}\,\left\|\theta_{0}-u\right\|_{2}^{2}}{\,(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right]$
iterations to ensure that $f(\theta_{T})-f(u)\leq\epsilon$.
Compared to Corollary 1, the above result only focuses on the case when $L_{0}=0,\,\omega=0$ and requires an additional assumption that $f(\theta)$ is $L$-smooth. These conditions are satisfied for both logistic regression and multi-class classification. We use the above result and prove the following corollary for logistic regression on separable data.
Corollary 2.
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon>0$, GD-LS with $\eta_{\max}=\infty$ requires
$T\geq\frac{6\,c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
to ensure that $f(\theta_{T})\leq 2\,\epsilon$.
Proof Sketch.
We use a proof technique similar to that in Ji & Telgarsky (2018), and consider the max-margin solution $u^{*}$ where $\left\|u^{*}\right\|=1$ and $\gamma=\min_{i}\,y_{i}\langle x_{i},u^{*}\rangle$. Note that for any scalar $\beta>0$, $f(\beta u^{*})\leq\exp(-\beta\gamma)$. Choosing $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$, setting the comparator $u=\beta\,u^{*}$ in Theorem 2 and using the bound on $f(\beta u^{*})$ completes the proof. ∎
Hence, on linearly separable data, GD-LS can achieve a linear rate of convergence for logistic regression (see Fig. 2 for an experimental validation), resulting in an exponential improvement over the standard $O(1/\epsilon)$ rate for smooth, convex functions. Finally, the result in Corollary 7 is better than the $O(1/\sqrt{\epsilon})$ rate for GD with large (than $1/L$) constant step-sizes (Wu et al., 2024).
Figure 2: Comparing GD-LS with $c=1/2$ and $\eta_{\max}=10^{8}$ and GD(1/L) for unregularized logistic regression on a synthetic separable dataset with $\gamma=0.1$, $n=10^{4}$ and $d=200$. (Left) Sub-optimality plot: GD-LS converges linearly, while GD(1/L) has a sublinear convergence. (Right) Step-size plot: The GD-LS step-size increases non-monotonically and becomes larger than $10^{5}$ in $40$ iterations.
An interesting question is whether the above linear convergence rate can only be achieved by methods such as GD-LS which ensure that the loss is monotonically decreasing. In Section C.1, we answer this in the negative, and show that the commonly used Polyak step-size (Polyak, 1987) can also achieve the fast convergence rate in Theorem 2 and consequently, result in linear convergence for logistic regression on separable data.
In the next section, we consider the convergence of GD-LS on non-convex losses.
5 GD-LS for Non-convex Losses
In this section, we consider non-convex losses that satisfy two alternative gradient domination conditions which enable convergence to the global optimum. In Section 5.1, we analyze the convergence of GD-LS for objectives corresponding to the softmax policy gradient in reinforcement learning. In Section 5.2, we consider the standard Polyak-Łojasiewicz (PL) condition (Karimi et al., 2016; Polyak, 1987) for generalized linear models with a logistic link.
Assumption 4.
$f$ satisfies a non-uniform gradient domination condition with constant $\zeta\geq 1$, if there exists a $\mu(\theta)>0$ s.t. for all $\theta$,
$\left\|\nabla f(\theta)\right\|^{\zeta}\geq\mu(\theta)\,[f(\theta)-f^{*}].$
Gradient domination or Łojasiewicz conditions are satisfied for matrix factorization (Ward & Kolda, 2023), policy gradient in reinforcement learning (Mei et al., 2020) and generalized linear models (Mei et al., 2021). This property has been exploited to prove global convergence guarantees for first-order methods (Karimi et al., 2016; Mei et al., 2021).
5.1 Gradient domination with $\zeta=1$
We use softmax policy optimization for multi-armed bandits (MAB) as a concrete example that satisfies Assumptions 1, 2 and 3 and Assumption 4 with $\zeta=1$. In particular, we consider an MAB problem with deterministic, known rewards that is often used as a testbed to evaluate policy gradient methods (Xiao, 2022; Mei et al., 2020; Lu et al., 2024). We prove the following proposition in Appendix A.
Proposition 3.
Given an MAB problem with $K$ arms and known deterministic rewards $r\in[0,1]^{K}$, consider the class of softmax policies $\pi_{\theta}\in\Delta_{K}$ parameterized by $\theta\in\mathbb{R}^{K}$ s.t. $\pi_{\theta}(a)=\frac{\exp(\theta(a))}{\sum_{a^{\prime}}\exp(\theta(a^{\prime}))}$. The loss corresponding to the bandit problem is given by:
$f(\theta)=r(a^{*})-\langle\pi_{\theta},r\rangle\,,$
where $a^{*}:=\operatorname*{arg\,max}_{a\in[K]}r(a)$ is the optimal arm. $f(\theta)$ is non-negative, satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\frac{3\sqrt{2}}{\Delta}$ssumption 3 with $\nu=\frac{\sqrt{2}}{\Delta}$ and $\omega=0$ and Assumption 4 with $\zeta=1$ and $\mu(\theta)=\pi_{\theta}(a^{*})$. Here, $\Delta:=\max_{a\neq a^{*}}r(a^{*})-r(a)$ is the reward gap and quantifies the problem difficulty.
Softmax policy gradient methods (Williams, 1992) optimize the above non-convex objective using gradient descent, and have been analyzed recently (Mei et al., 2020; Agarwal et al., 2021). We aim to use GD-LS to optimize the objective defined in Proposition 3 for which the GD update is given by: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla f(\theta_{t})=\theta_{t}+{\eta_{t}}\nabla_{\theta}\langle\pi_{\theta},r\rangle$, and the Armijo condition is equal to:
$\langle\pi_{\theta_{t+1}},r\rangle$ $\geq\langle\pi_{\theta},r\rangle+c{\eta_{t}}\,\left\|\nabla_{\theta}\langle\pi_{\theta},r\rangle\right\|_{2}^{2}$
Hence, unlike in Lu et al. (2024), implementing GD-LS does not require knowledge of $r(a^{*})$, and hence the identity of the optimal arm. In Proposition 6 in Appendix A, we show that, under additional assumptions, the softmax policy gradient objective for tabular Markov decision processes (MDPs) also satisfies the Assumptions 1, 2, 3 and 4 with $L_{0}=0$ and $\omega=0$. Consequently, we characterize the convergence rate of GD-LS on such problems.
Corollary 3.
For a fixed $\epsilon>0$, assuming $f(\theta)$ satisfies Assumptions 1, 2 and 3 with $L_{0}=0,\,\omega=0$ and Assumption 4 with $\zeta=1$, GD-LS with $\eta_{\max}=\infty$, requires
$T$ $\geq\max\left\{1,\frac{2\,\lambda_{1}}{\mu^{2}}\right\}\,\left(\frac{{f^{*}}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-{f^{*}}}{\epsilon}\right)$
iterations to ensure $f(\theta_{T})\leq\epsilon$ where $\mu:=\min_{t\in[T]}\mu(\theta_{t})$.
In order to better understand the implications of Corollary 3, we instantiate the above result for MAB and obtain the following corollary.
Corollary 4.
For an MAB problem with $K$ arms, rewards bounded in $[0,1]$ and reward gap equal to $\Delta$, GD-LS with a uniform initialization i.e. $\pi_{\theta_{0}}(a)=\frac{1}{K}$ for all $a$, $c=\frac{1}{2}$, $\eta_{\max}=\infty$ requires $T\geq\frac{12^{2}\,K^{2}}{\Delta}\,\ln\left(\frac{1}{\epsilon}\right)$ iterations to guarantee $\langle\pi_{\theta_{T}},r\rangle\geq r(a^{*})-\epsilon$.
Hence, for MAB problems, GD-LS converges at a linear rate. Under additional assumptions, a similar result also holds for tabular MDPs. This is in contrast to GD(1/L) which can only attain an $\Omega\left(\frac{1}{\epsilon}\right)$ convergence rate for both bandits and MDPs (Mei et al., 2020, Theorem 9, 10). The convergence rate of GD-LS matches that of algorithms explicitly designed for this problem, including GD with a specific line-search that requires knowledge of $r(a^{*})$ (Lu et al., 2024), GD with specific increasing step-sizes (Liu et al., 2024), normalized GD (Mei et al., 2021), natural policy gradient (Xiao, 2022) and mirror descent with a log-sum-exp mirror map (Asad et al., 2024).
We note that Assumptions 1, 2 and 3 and Assumption 4 with $\zeta=1$ and $f^{*}=0$ are also satisfied when using the (i) exponential loss to train (ii) two-layer neural networks with a smoothed leaky-ReLU non-linearity and (iii) assuming that the training data is linearly separable (Taheri & Thrampoulidis, 2023, Lemmas 3,5). In this setting, Theorem 1 in Taheri & Thrampoulidis (2023) shows that normalized GD can result in linear convergence for the resulting non-convex objective. Since this problem also satisfies the required assumptions for Corollary 3, we conclude that GD-LS also converges linearly and unlike normalized GD, it does not require knowing problem-dependent constants. Hence, our results demonstrate the universality of GD-LS.
5.2 Gradient domination with $\zeta=2$
We use generalized linear models (GLM) with a logistic link function as an example. In Proposition 2, we have seen that the objective satisfies Assumptions 1, 2 and 3 with non-zero $L_{0},L_{1}$, $\nu,\omega$. It also satisfies Assumption 4 with $\zeta=2$.
Lemma 2 (Lemma 9 in (Mei et al., 2021)).
If $\sigma(\cdot)$ is the sigmoid function and $\pi_{i}(\theta):=\sigma(\langle x_{i},\theta\rangle)$, assuming that for all $i\in[n]$, $\left\|x\right\|_{i}\leq 1$, $y_{i}=\pi_{i}(\theta^{*})$ such that $\left\|\theta^{*}\right\|\leq D<\infty$ and $\upsilon(\theta):=\min_{i\in[n]}{\left\{\pi_{i}(\theta)\cdot\left(1-\pi_{i}(\theta)\right)\right\}}$, then the GLM objective in Eq. 4 satisfies Assumption 4 with $\zeta=2$ and $\mu(\theta)=64\,[\upsilon(\theta)]^{2}\,[\min\{\upsilon(\theta),\upsilon(\theta^{*})\}]^{2}$.
Similar to logistic regression, the PL constant $\mu$ depends on $\upsilon(\theta)$. However, unlike logistic regression where $y_{i}\in\{0,1\}$ and $\left\|\theta^{*}\right\|$ can be infinite on separable data, for GLMs, $y_{i}\in(0,1)$ and $\left\|\theta^{*}\right\|$ is bounded. Consequently, $\mu(\theta^{*})>0$ and as long as $\left\|\theta_{t}\right\|<\infty$ for any iterate $t\in[T]$, $\mu(\theta)$ is bounded away from zero. However, we note that this does not preclude the case where the iterates initially diverge away from the solution, resulting in large $\left\|\theta_{t}\right\|$ and small (but non-zero) $\mu(\theta)$. Furthermore, the realizability condition in Lemma 2 implies that $f^{*}=0$. Given these considerations, we prove the following theorem in Appendix D.
Theorem 3.
For a fixed $\epsilon\in\left(0,\frac{\lambda_{0}}{\lambda_{1}}\right)$, if $f$ satisfies Assumptions 1, 2 and 3 and Assumption 4 with $\zeta=2$ with $f^{*}=0$ and if $\mu:=\min_{t\in[T]}\mu(\theta_{t})$, GD-LS with $\eta_{\max}=\infty$, requires
$T\geq\frac{2}{\mu}\,\left[\lambda_{1}{f}(\theta_{0})+\lambda_{0}\,\ln\left(\frac{\lambda_{0}}{\lambda_{1}\,\epsilon}\right)\right]$
iterations to ensure that $f(\theta_{T})\leq\epsilon$.
Proof Sketch.
Using the gradient domination condition with the Armijo condition in Eq. 6 and the lower-bound on ${\eta_{t}}$ from Lemma 1, we get that, $f(\theta_{t+1})\leq f(\theta_{t})-\frac{\mu\,f(\theta_{t})}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$. For the first $\tau$ iterations, we have that $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$. Since $\lambda_{0}\leq\lambda_{1}\,f(\theta_{t})$, by recursing for $\tau$ iterations, we conclude that $f(\theta_{\tau})\leq f(\theta_{0})-\tau\,\frac{\mu}{2\,\lambda_{1}}$, meaning that $\tau\leq\frac{2\lambda_{1}\,f(\theta_{0})}{\mu}$. This corresponds to Phase 1. that terminates in $O(1)$ iterations. For all iterations $t\in[\tau,T]$, $f(\theta_{t})\leq\frac{\lambda_{0}}{\lambda_{1}}$ and hence $\lambda_{0}\geq\lambda_{1}\,f(\theta_{t})$. Using this relation and following the standard proof for uniformly smooth functions satisfying the PL inequality completes the proof for Phase 2. which terminates in $O(\ln(1/\epsilon))$ iterations. Putting together the results for both phases completes the proof. ∎
The above result shows that GD-LS converges linearly, where the convergence rate depends on the ratio $\nicefrac{{\lambda_{0}}}{{\lambda_{1}}}$. On the other hand, for an $L$ uniformly-smooth function satisfying the PL condition, GD(1/L) requires $O\left(\frac{L}{\mu}\,\ln\left(\frac{f(\theta_{0})}{\epsilon}\right)\right)$ iterations (Karimi et al., 2016). Ignoring the constant first term independent of $\epsilon$ and assuming $\lambda_{0}\approx L$, we can see that the result in Theorem 3 is better than the standard rate when $\nicefrac{{\lambda_{0}}}{{\lambda_{1}}}$ is smaller than $f(\theta_{0})$. It is important to note that since GD-LS can automatically (without any change in the algorithm) exploit the uniform smoothness as well and obtain the standard result, the number of iterations required for GD-LS is $\min\left\{O\left(\frac{L}{\mu}\,\ln\left(\frac{f(\theta_{0})}{\epsilon}\right)\right),O\left(\frac{\lambda_{0}}{\mu}\,\ln\left(\frac{\lambda_{0}}{\lambda_{1}\,\epsilon}\right)\right)\right\}$, meaning that GD-LS should always converge at least as fast as GD(1/L). Since the GLM objective is also uniformly smooth (Mei et al., 2021, Lemma 10), we empirically verify our hypothesis in Fig. 3.
Figure 3: Comparing GD-LS with $c=1/2$, $\eta_{\max}=10^{4}$ and GD(1/L) for GLMs on a synthetic dataset with $n=10^{4}$, $d=200$, $\left\|\theta^{*}\right\|=1$. GD-LS converges faster than GD(1/L), and demonstrates Phase 1 behaviour initially, followed by a linear rate in Phase 2.
Comparing Theorem 3 for GLMs to the existing results, we note that Hazan et al. (2015, Lemma 3.1) show that the GLM objective is locally quasi-convex, and use this property to derive a slower $O(1/\epsilon^{2})$ convergence rate for normalized GD with a decreasing step-size (Hazan et al., 2015, Theorem 4.1). Finally, for the GLM objective, Mei et al. (2021) propose a novel variant of normalized GD and prove that it converges linearly, with a better constant dependence than GD. However, their method requires the knowledge of $\mu$, making it difficult to implement. On the other hand, GD-LS does not require parameter tuning and achieves similar theoretical results as these specialized methods.
In the next section, we show the benefits of using a line-search in the stochastic setting.
6 SGD with Stochastic Line-search
In this section, we analyze the convergence of stochastic gradient descent (SGD) (Robbins & Monro, 1951) with a stochastic variant of the Armijo line-search (referred to as SGD-SLS) proposed in Vaswani et al. (2019b). We focus on the convex, finite-sum setting, and consider minimizing $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$ where each $f_{i}$ is convex and satisfies Assumptions 1, 2 and 3. For ease of exposition, we assume that each $f_{i}$ is $(L_{0},L_{1})$ non-uniform smooth, and note that it is straightforward to analyze the case where the $f_{i}$’s have different smoothness constants. Binary classification using logistic regression and multi-class classification using the cross-entropy loss are examples of such an objective.
At iteration $t\in[T]$, SGD randomly samples a function ${f_{t}}$ from the $n$ functions in the finite-sum, computes its gradient and takes a descent step. Specifically,
$\theta_{t+1}=\theta_{t}-{\eta_{t}}\nabla{f_{t}}(\theta_{t})$ (8)
where $\nabla{f_{t}}(\theta_{t})$ is the gradient of the loss function chosen at iteration $t$. Each stochastic gradient $\nabla{f_{t}}(\theta_{t})$ is unbiased, implying that $\mathbb{E}_{t}\left[\nabla f_{t}(\theta)\right]=\nabla f(\theta)$. In order to estimate ${\eta_{t}}$, SGD-SLS uses the stochastic analog of the Armijo condition in Eq. 6. In particular, starting from $\eta_{\max}$, SLS uses a backtracking procedure and returns the largest step-size ${\eta_{t}}$ that satisfies: ${\eta_{t}}\leq\eta_{\max}$ and,
${f_{t}}(\theta_{t}-{\eta_{t}}\nabla{f_{t}}(\theta_{t}))$ $\leq{f_{t}}(\theta_{t})-c\,{\eta_{t}}\left\|\nabla{f_{t}}(\theta_{t})\right\|_{2}^{2}.$ (9)
Note that the stochastic Armijo condition only involves the sampled function and its gradient.
In order to analyze the convergence of SGD-SLS, we define $f_{i}^{*}:=\min f_{i}(\theta)$ as the minimum of function $i$ in the finite-sum and $\chi^{2}(\theta^{*}):=\mathbb{E}_{i}[f_{i}(\theta^{*})-f_{i}^{*}]$ as the “noise” in the stochastic gradients at the optimum (Loizou et al., 2021). In particular, if $\chi^{2}=0$, then each $f_{i}$ is minimized at $\theta^{*}$ implying that $\nabla f_{i}(\theta^{*})=0$. This special case is referred to as the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt & Roux, 2013) and is useful in practical machine learning; for example, it is approximately satisfied by over-parameterized neural networks (Zhang et al., 2017) or non-parametric regression (Liang & Rakhlin, 2018; Belkin et al., 2019). Furthermore, logistic regression on linearly separable data is an example of a smooth convex loss that satisfies the interpolation condition and is the main motivation for the subsequent analysis.
When minimizing uniformly-smooth convex functions in the interpolation setting, Vaswani et al. (2019a) proved that SGD-SLS converges to the optimum at the $O(1/\epsilon)$ rate matching GD and faster than the standard $O(1/\epsilon^{2})$ rate (Bottou et al., 2018) for SGD. Moreover, SGD-SLS does not require the knowledge of the global smoothness constant, making it an attractive choice in practice. Motivated by our linear convergence results in Section 4, we analyze the convergence of SGD-SLS for convex functions that satisfy Assumptions 1, 2 and 3.
Since SLS involves an Armijo line-search for one (randomly chosen) function in each iteration, we can follow the same argument as in Lemma 1 and show that the step-size in each iteration is lower-bounded as ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{1}{\lambda_{0}\,+\lambda_{1}\,{f_{t}}(\theta_{t})}\right\}$. Given this result, we first prove the following lemma in Appendix E.
Lemma 3.
For a fixed $\epsilon>0$, assuming $f(\theta):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$ where each $f_{i}$ is convex, $L$ uniform smooth, satisfies Assumptions 1, 2 and 3 with $L_{0}=0,\,\omega=0$, if $\Delta_{t}:=\mathbb{E}[\left\|\theta_{t}-u\right\|_{2}^{2}]$, then, SGD-SLS guarantees that:
$\Delta_{T}$ $\leq\Delta_{0}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]$
$+2\eta_{\max}\,\chi^{2}(u)\,T$
where $u$ is an arbitrary comparator s.t. $f(u)<\mathbb{E}[f(\theta_{T})]$, $C:=\frac{(2c-1)}{c\,\lambda_{1}}$ and $\chi^{2}(u):=\mathbb{E}_{i}[f_{i}(u)-\min f_{i}(\theta)]$ is the noise in the stochastic gradients at $u$.
Since logistic regression on linearly separable data satisfies Assumptions 1, 2 and 3 and the interpolation condition, we follow a similar strategy as in Corollary 7 and prove the following corollary for logistic regression on separable data.
Corollary 5.
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|=1$, for a fixed $\epsilon\in\left(0,\frac{1}{8}\right)$, SGD-SLS with $\eta_{\max}=\frac{C}{\epsilon}$ where $C=\frac{(1-c)\,(2c-1)}{c\,\lambda_{1}}$ requires $T$ iterations to ensure that $\mathbb{E}[f(\theta_{T})]\leq 2\epsilon$ where,
$T\geq\frac{6\,c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}.$
Hence, SGD-SLS requires $O\left(\ln\left(\frac{1}{\epsilon}\right)^{2}\right)$ iterations (and hence gradient evaluations) to ensure convergence to an $\epsilon$ neighbourhood. This is an exponential improvement over the standard $O(1/\epsilon^{2})$ convergence rate for SGD on convex, uniformly smooth functions, where the improvements stem from effectively exploiting both interpolation and non-uniform smoothness.
7 Conclusion
We explored the theoretical properties of GD-LS for a class of functions satisfying non-uniform smoothness. For a range of practically useful convex and non-convex functions, we proved that Armijo-LS can effectively adapt to the objective’s structural properties and enable faster convergence for GD. In particular, we showed that GD-LS can either match or provably improve upon the sublinear rate of GD(1/L), and do so without relying on the knowledge of problem-dependent constants. Furthermore, for specific problems in supervised learning and reinforcement learning, we demonstrated that GD-LS can match the fast convergence of algorithms tailored for these problems. In conclusion, our results show the universality and effectiveness of GD-LS. We believe that investigating the behaviour of GD-LS for a broader class of non-convex functions, and formally characterizing the advantage of Armijo-LS for other algorithms (such as Nesterov accelerated gradient) remain important directions for future work.
Acknowledgments
We thank Damien Scieur, Mark Schmidt, Curtis Fox, Michael Lu and Siyi Meng for helpful discussions and feedback, and Anh Dang for help with the initial experiments.
Supplementary Material
Organization of the Appendix
Appendix A Proofs for Section 2
In order to prove that commonly used functions in machine learning satisfy the assumptions in Section 2, we will require some additional assumptions and intermediate results.
Assumption 5.
$f$ is twice-differentiable and satisfies the following non-uniform smoothness property: for constants $L_{c}$ , $L_{g}>0$,
$\left\|\nabla^{2}f(\theta)\right\|\leq L_{c}+L_{g}\,\left\|\nabla f(\theta)\right\|$
Unlike Assumption 2Assumption 5 corresponds to the standard non-uniform smoothness assumption made in the literature (Zhang et al., 2019, 2020) and implies the following result.
Lemma 4 (Lemma A3 (Zhang et al., 2020)).
If $f$ satisfies Assumption 5, then the following inequality holds for all $x,y$ such that $\left\|x-y\right\|\leq\frac{q}{L_{g}}$ where $q>0$ is a constant,
$f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{AL_{c}+B\,L_{g}\,\left\|\nabla f(x)\right\|}{2}\left\|y-x\right\|_{2}^{2}\,,$
where $A:=1+e^{q}-\frac{e^{q}-1}{q}$ and $B:=\frac{e^{q}-1}{q}$.
Lemma 5.
For a finite-sum objective, $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$, if, for all $i$, $f_{i}$ satisfies Assumption 3 with the constants equal to $\nu,\omega$, then,
$\left\|\nabla f(\theta)\right\|\leq\nu\,f(\theta)+\omega$
Furthermore, if for all $i$, $f_{i}$ also satisfies Assumption 5 with constants $L_{c},L_{g}$, then, the following inequalities hold $L_{0}:=L_{c}+L_{g}\,\omega$ and $L_{1}:=\nu\,L_{g}$:
(a) for all $\theta$,
$\left\|\nabla^{2}f(\theta)\right\|\leq L_{0}+L_{1}\,f(\theta)$
(b) for all $x,y$ such that $\left\|x-y\right\|\leq\frac{q}{L_{1}}$ where $q\geq 1$ is a constant,
$f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{AL_{0}+B\,L_{1}\,f(x)}{2}\,\left\|y-x\right\|_{2}^{2}\,,$
where $A:=1+e^{q}-\frac{e^{q}-1}{q}$ and $B:=\frac{e^{q}-1}{q}$.
Hence, if $f_{i}(\theta)$ satisfies Assumption 3 and Assumption 5, then $f(\theta)$ satisfies Assumption 2 and Assumption 3.
Proof.
For the first part, note that,
$\left\|\nabla f(\theta)\right\|$ $=\left\|\frac{1}{n}\,\sum_{i}\nabla f_{i}(\theta)\right\|\leq\frac{1}{n}\,\sum_{i}\left\|\nabla f_{i}(\theta)\right\|\leq\frac{\nu}{n}\,\sum_{i}f_{i}(\theta)+\omega=\nu\,f(\theta)+\omega\,.$
For the second part,
$\left\|\nabla^{2}f(\theta)\right\|$ $=\left\|\frac{1}{n}\,\sum_{i}\nabla^{2}f_{i}(\theta)\right\|\leq\frac{1}{n}\,\sum_{i}\left\|\nabla^{2}f_{i}(\theta)\right\|\leq L_{c}+\frac{L_{g}}{n}\,\sum_{i}\left\|\nabla f_{i}(\theta)\right\|$
$\leq L_{c}+\frac{L_{g}}{n}\,\sum_{i}[\nu\,f_{i}(\theta)+\omega]=[L_{c}+L_{g}\,\omega]+\nu\,L_{g}\,f(\theta)$
For the third part, using Lemma 5, for all $i$,
$f_{i}(y)$ $\leq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{AL_{c}+B\,L_{g}\,\left\|\nabla f_{i}(x)\right\|}{2}\left\|y-x\right\|_{2}^{2}$
$\leq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{AL_{c}+B\,L_{g}\,[\nu\,f_{i}(x)+\omega]}{2}\left\|y-x\right\|_{2}^{2}$ (Since $f_{i}$ satisfies Assumption 3 with constants $\nu,\omega$)
$=f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{A\left(L_{c}+\frac{B}{A}\,L_{g}\,\omega\right)+B\,L_{g}\,\nu\,f_{i}(x)}{2}\left\|y-x\right\|_{2}^{2}$
$\leq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{A\left(L_{c}+L_{g}\,\omega\right)+B\,L_{g}\,\nu\,f_{i}(x)}{2}\left\|y-x\right\|_{2}^{2}$ (Since $B\leq A$)
Summing the LHS and RHS for $i=1$ to $n$ and dividing by $n$ completes the proof. ∎
Hence, if the conditions of Lemma 5 are satisfied, then the non-uniform smoothness condition in Assumption 2 is satisfied. The following propositions show that such a non-uniform smoothness condition is satisfied for linear logistic regression, linear model with an exponential loss, linear multi-class classification using the cross-entropy loss, generalized linear models with a logistic link function and the softmax policy gradient objective for multi-armed bandits and tabular MDPs.
See 1
Proof.
Clearly, $f_{i}(\theta)\geq 0$ and hence $f(\theta)\geq 0$ for all $\theta$. $f(\theta)$ is a finite-sum objective. Calculating the gradient and hessian for $f_{i}(\theta):=\ln(1+\exp(-y_{i}\langle x_{i},\theta\rangle)$,
$\nabla f_{i}(\theta)$ $=\frac{-\exp(-y_{i}\langle x_{i},\theta\rangle)}{1+\exp(-y_{i}\langle\,x_{i},\theta\rangle)}y_{i}\,x_{i}\quad\text{;}\quad\nabla^{2}f_{i}(\theta)=\frac{1}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}\,\frac{\exp(-y_{i}\,\langle x_{i},\theta\rangle)}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}\,y_{i}^{2}\,x_{i}\,x_{i}^{T}$
$\left\|\nabla f_{i}(\theta)\right\|$ $=\frac{\exp(-y_{i}\langle x_{i},\theta\rangle)}{1+\exp(-y_{i}\langle\,x_{i},\theta\rangle)}\left\|x_{i}\right\|$
$\implies\left\|\nabla f_{i}(\theta)\right\|$ $\leq\left\|x_{i}\right\|\,\ln\left(1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)\right)=\left\|x_{i}\right\|\,f_{i}(\theta)$ (For all $x$, $\frac{x}{1+x}\leq\ln(1+x)$)
Hence, for all $i$, $f_{i}$ satisfies Assumption 3 with $\nu=\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. Bounding the Hessian,
$\nabla^{2}f_{i}(\theta)$ $\preceq\frac{\exp(-y_{i}\,\langle x_{i},\theta\rangle)}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}\,y_{i}^{2}x_{i}x_{i}^{T}=\frac{\exp(-y_{i}\,\langle x_{i},\theta\rangle)}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}\,x_{i}x_{i}^{T}$ (For all $x$, $\frac{1}{1+e^{x}}\leq 1$ and $y_{i}^{2}=1$)
$\implies\left\|\nabla^{2}f_{i}(\theta)\right\|$ $\leq\frac{\exp(-y_{i}\,\langle x_{i},\theta\rangle)}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}\,\left\|x_{i}\right\|_{2}^{2}=\left\|x_{i}\right\|\,\left\|\nabla f_{i}(\theta)\right\|$
Hence, for all $i$, $f_{i}$ satisfies Assumption 5 with $L_{c}=0$ and $L_{g}=\max_{i}\left\|x_{i}\right\|$. Using Lemma 5, we conclude that $f(\theta)$ satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=L_{g}\,\nu=\max_{i}\left\|x_{i}\right\|_{2}^{2}$, and Assumption 3 with $\nu=\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. ∎
See 2
Proof.
Clearly, $f_{i}(\theta)\geq 0$ and hence $f(\theta)\geq 0$ for all $\theta$. $f(\theta)$ is a finite-sum objective. Calculating the gradient and hessian for $f_{i}(\theta)=\frac{1}{2}\,\left(\pi_{i}(\theta)-y_{i}\right)^{2}$,
$\nabla f_{i}(\theta)$ $=\left(\pi_{i}(\theta)-y_{i}\right)\,\frac{1}{1+\exp(-\langle x_{i},\theta\rangle)}\,\frac{\exp(-\langle x_{i},\theta\rangle)}{1+\exp(-\langle x_{i},\theta\rangle)}\,x_{i}$
$\implies\left\|\nabla f_{i}(\theta)\right\|$ $=|\pi_{i}(\theta)-y_{i}|\,\pi_{i}(\theta)\,\underbrace{(1-\pi_{i}(\theta))}_{\leq 1}\,\left\|x_{i}\right\|\leq|\pi_{i}(\theta)-y_{i}|\,|\pi_{i}(\theta)-y_{i}+y_{i}|\,\left\|x_{i}\right\|$
$\leq\left[|\pi_{i}(\theta)-y_{i}|\,|\pi_{i}(\theta)-y_{i}|+y_{i}\,|\pi_{i}(\theta)-y_{i}|\right]\,\left\|x_{i}\right\|$ (Triangle inequality)
$=2\,\left\|x_{i}\right\|\,\left[\frac{1}{2}\,\left(\pi_{i}(\theta)-y_{i}\right)^{2}\right]+\left\|x_{i}\right\|$ (Since $y_{i}\,|\pi_{i}(\theta)-y_{i}|\in[0,1]$)
$\implies\left\|\nabla f_{i}(\theta)\right\|$ $\leq 2\,\left\|x_{i}\right\|\,f_{i}(\theta)+\left\|x_{i}\right\|$
Hence, for all $i$, $f_{i}$ satisfies Assumption 3 with $\nu=2\,\max_{i\in[n]}\left\|x_{i}\right\|$ and $\omega=\max_{i\in[n]}\left\|x_{i}\right\|$. Calculating the Hessian,
$\nabla^{2}f_{i}(\theta)$ $=[1-2\,\pi_{i}(\theta)]\,\pi_{i}(\theta)\,[1-\pi_{i}(\theta)]\,[\pi_{i}(\theta)-y_{i}]\,x_{i}\,x_{i}^{T}+[\pi_{i}(\theta)]^{2}\,[1-\pi_{i}(\theta)]^{2}\,x_{i}\,x_{i}^{T}$
$\implies\left\|\nabla^{2}f_{i}(\theta)\right\|$ $=\left[\underbrace{|1-2\,\pi_{i}(\theta)|}_{\leq 1}\,\underbrace{\pi_{i}(\theta)\,[1-\pi_{i}(\theta)]\,|\pi_{i}(\theta)-y_{i}|\left\|x_{i}\right\|}_{=\left\|\nabla f_{i}(\theta)\right\|}+\underbrace{[\pi_{i}(\theta)]^{2}\,[1-\pi_{i}(\theta)]^{2}}_{\leq\frac{1}{16}}\left\|x_{i}\right\|\right]\,\left\|x_{i}\right\|$ (Triangle Inequality)
$\implies\left\|\nabla^{2}f_{i}(\theta)\right\|$ $\leq\left\|x_{i}\right\|\,\left\|\nabla f_{i}(\theta)\right\|+\frac{1}{16}\left\|x_{i}\right\|_{2}^{2}$
Hence, for all $i$, $f_{i}(\theta)$ satisfies Assumption 5 with $L_{c}=\frac{1}{16}\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and $L_{g}=\max_{i\in[n]}\left\|x_{i}\right\|$. Using Lemma 5, we conclude that $f(\theta)$ satisfies Assumption 2 with $L_{0}=\frac{17}{16}\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and $L_{1}=2\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$, and Assumption 3 with $\nu=2\max_{i}\left\|x_{i}\right\|$ and $\omega=\max_{i}\left\|x_{i}\right\|$. ∎
Proposition 4.
Consider $n$ points where $x_{i}\in\mathbb{R}^{d}$ are the features and $y_{i}\in\{0,1\}^{C}$ are the corresponding one-hot label vectors for $C$ classes. Multi-class classification with the cross-entropy objective is given as:
$f(\theta)$ $=\frac{1}{n}\sum_{m=1}^{n}\text{KL}(y^{m}||\pi^{m}_{\theta})\,,\text{ where $\forall m\in[n]$, $\pi^{m}_{\theta}\in\Delta_{C}$ s.t. $\forall i\in[C]$, }\pi^{m}_{\theta}(i)=\frac{\exp(\langle x_{m},\theta_{i}\rangle)}{\sum_{k=1}^{C}\exp(\langle x_{m},\theta_{k}\rangle)}\,,$
where $\theta_{i}\in\mathbb{R}^{d}$ for $i\in[C]$ and $\theta=[\theta_{1},\theta_{2},\ldots,\theta_{C}]$. Multi-class logistic regression satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=4\,\max_{m\in[n]}\left\|x\right\|^{2}_{1}$, and Assumption 3 with $\nu=2\max_{i}\left\|x_{i}\right\|$ and $\omega=0$.
Proof.
Let us consider a single input-output pair $(x,y)$ and calculate the gradient for a single function in the finite-sum. Define $\ell(\theta):=\text{KL}(y||\pi_{\theta})$ where $y$ is a $C$-dimensional one-hot vector, $x\in\mathbb{R}^{d}$ and $\pi_{\theta}\in\Delta_{C}$ s.t. $\pi_{\theta}(i)=\frac{\exp(\langle x,\theta_{i}\rangle)}{\sum_{k=1}^{C}\exp(\langle x,\theta_{k}\rangle)}$
$\frac{\partial\ell(\theta)}{\partial\theta_{i}}$ $=[\pi_{\theta}(i)-y_{i}]\,x\implies\left\|\frac{\partial\ell(\theta)}{\partial\theta_{i}}\right\|_{1}=|\pi_{\theta}(i)-y_{i}|\,\left\|x\right\|_{1}$
$\implies\left\|\nabla_{\theta}\ell(\theta)\right\|_{1}$ $=\left\|x\right\|_{1}\,\sum_{i=1}^{C}|\pi_{\theta}(i)-y_{i}|$
Since $y$ is a one-hot vector, let $i^{*}$ be the index corresponding to the non-zero entry. Hence, $y_{i^{*}}=1$ and for all $j\neq i^{*}$, $y_{j}=0$. With this,
$\leq\left\|x\right\|_{1}\,\sum_{i\neq i^{*}}\pi_{\theta}(i)+[1-\pi_{\theta}(i^{*})]=2\,[1-\pi_{\theta}(i^{*})]\,\left\|x\right\|_{1}$
$\leq 2\,\left\|x\right\|_{1}\,\ln\left(\frac{1}{\pi_{\theta}(i^{*})}\right)$ (For all $z\in[0,1]$, $1-z\leq\ln(1/z)$)
$=2\,\left\|x\right\|_{1}\,\sum_{i=1}^{C}y_{i}\,\ln\left(\frac{y_{i}}{\pi_{\theta}(i)}\right)$ (Using that $y$ is a one-hot vector)
$\implies\left\|\nabla_{\theta}\ell(\theta)\right\|_{1}$ $\leq 2\,\left\|x\right\|_{1}\,\text{KL}(y||\pi_{\theta})=2\,\left\|x\right\|_{1}\,\ell(\theta)\,.$
Hence, $\ell(\theta)$ satisfies Assumption 3 with $\nu=2\left\|x\right\|_{1}$ and $\omega=0$. Let us now bound the Hessian. The Hessian can be written as a Kronecker product of a $C\times C$ matrix which corresponds to the Jacobian of the softmax function, and a $d\times d$ rank-one matrix formed using the features. Specifically,
$\nabla^{2}\ell(\theta)$ $=\underbrace{H}_{C\times C}\,\underbrace{xx^{T}}_{d\times d}\text{ where, }H:=\text{diag}(\pi_{\theta})-\pi_{\theta}\,\pi_{\theta}^{T}$
$\implies\left\|\nabla^{2}\ell(\theta)\right\|$ $\leq\left\|x\right\|_{2}^{2}\,\left\|H\right\|$
Since $H$ is a square symmetric PSD matrix, $\left\|H\right\|=\lambda_{\max}[H]$. By the Gershgorin circle theorem, $\lambda_{\max}[H]\leq\max_{i}\sum_{j=1}^{C}|H_{i,j}|$. Calculating the row sums, we conclude that $\left\|H\right\|\leq\lambda_{\max}[H]\leq 2\max_{i}\pi_{\theta}(i)\,(1-\pi_{\theta}(i))$. Hence,
$\left\|\nabla^{2}\ell(\theta)\right\|$ $\leq 2\,\left\|x\right\|_{2}^{2}\,\max_{i}\pi_{\theta}(i)\,(1-\pi_{\theta}(i))\leq 2\,\left\|x\right\|\,\frac{\max_{i}\pi_{\theta}(i)\,(1-\pi_{\theta}(i))}{\sum_{i=1}^{C}|\pi_{\theta}(i)-y_{i}|}\,\left\|\nabla\ell(\theta)\right\|_{1}$
Let $j^{*}:=\operatorname*{arg\,max}\pi_{\theta}(i)\,(1-\pi_{\theta}(i))$. Using that $\sum_{i=1}^{C}|\pi_{\theta}(i)-y_{i}|\geq|\pi_{\theta}(j^{*})-y_{j^{*}}|$ and $\left\|x\right\|_{2}\leq\left\|x\right\|_{1}$,
$\leq 2\,\left\|x\right\|_{1}\,\frac{\pi_{\theta}(j^{*})\,(1-\pi_{\theta}(j^{*}))}{|\pi_{\theta}(j^{*})-y_{j^{*}}|}\,\left\|\nabla\ell(\theta)\right\|_{1}$
$\implies\left\|\nabla^{2}\ell(\theta)\right\|$ $\leq 2\left\|x\right\|_{1}\,\left\|\nabla\ell(\theta)\right\|_{1}$ (Since $y_{j^{*}}\in\{0,1\}$ and $\pi_{\theta}(j^{*})\in[0,1]$)
Hence, for a single $(x,y)$ pair, we can conclude that, $\ell(\theta)$ satisfies Assumption 5 with $L_{g}=2\left\|x\right\|_{1}$.
Combining the above results implies that for all $m\in[n]$, $\text{KL}(y^{m}||\pi^{m}_{\theta})$ satisfies  Assumption 3 with $\nu=2\max_{m\in[n]}\left\|x\right\|_{1}$ and Assumption 5 with $L_{g}=2\max_{m\in[n]}\left\|x\right\|_{1}$. Since $f$ is a finite-sum, we use Lemma 5 to conclude that $f(\theta)$ satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=4\,\max_{m\in[n]}\left\|x\right\|^{2}_{1}$, and Assumption 3 with $\nu=2\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. ∎
Proposition 5.
Consider $n$ points where $x_{i}\in\mathbb{R}^{d}$ are the features and $y_{i}\in\{0,1\}$ are the corresponding labels. Binary classification with an exponential loss with the objective
$f(\theta):=\frac{1}{n}\,\sum_{i=1}^{n}\exp(-y_{i}\langle x_{i},\theta\rangle)\,,$
satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$, and Assumption 3 with $\nu=\max_{i}\left\|x_{i}\right\|$ and $\omega=0$.
Proof.
Clearly, $f_{i}(\theta)\geq 0$ and hence $f(\theta)\geq 0$ for all $\theta$. $f(\theta)$ is a finite-sum objective. Calculating the gradient and hessian for $f_{i}(\theta):=\exp(-y_{i}\langle x_{i},\theta\rangle)$,
$\nabla f_{i}(\theta)$ $=-\exp(-y_{i}\langle x_{i},\theta\rangle)y_{i}\,x_{i}\quad\text{;}\quad\nabla^{2}f_{i}(\theta)=\exp(-y_{i}\langle x_{i},\theta\rangle)\,y_{i}^{2}\,x_{i}\,x_{i}^{T}$
$\implies\left\|\nabla f_{i}(\theta)\right\|$ $=\exp(-y_{i}\langle x_{i},\theta\rangle)\,\left\|x_{i}\right\|=f_{i}(\theta)\,\left\|x_{i}\right\|$
Hence, for all $i$, $f_{i}$ satisfies Assumption 3 with $\nu=\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. Bounding the Hessian,
$\nabla^{2}f_{i}(\theta)$ $=\exp(-y_{i}\langle x_{i},\theta\rangle)\,y_{i}^{2}\,x_{i}\,x_{i}^{T}=\exp(-y_{i}\langle x_{i},\theta\rangle)\,x_{i}\,x_{i}^{T}$ ($y_{i}^{2}=1$)
$\implies\left\|\nabla^{2}f_{i}(\theta)\right\|$ $\leq\left\|x_{i}\right\|\,\left\|\nabla f_{i}(\theta)\right\|$
Hence, for all $i$, $f_{i}$ satisfies Assumption 5 with $L_{0}=0$ and $L_{g}=\max_{i}\left\|x_{i}\right\|$. Since $f$ is a finite-sum objective, using Lemma 5, we conclude that $f(\theta)$ satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=L_{g}\,\nu=\max_{i}\left\|x_{i}\right\|_{2}^{2}$. ∎
See 3
Proof.
From Mei et al. (2020, Lemma 17), we know that, if $\Delta:=\min_{a\neq a^{*}}r(a^{*})-r(a)$ is the minimum reward gap, then,
$\left\|\nabla\ell(\theta)\right\|=\left\|\nabla_{\theta}\langle\pi_{\theta},r\rangle\right\|$ $\leq\frac{\sqrt{2}}{\Delta}\,\ell(\theta)$
Hence, the loss for the bandit problem satisfies Assumption 3 with $\nu=\frac{\sqrt{2}}{\Delta}$ and $\omega=0$. From Mei et al. (2021, Lemma 2), we know that
$\left\|\nabla^{2}\ell(\theta)\right\|$ $=\left\|\nabla^{2}\langle\pi_{\theta},r\rangle\right\|\leq 3\,\left\|\nabla\langle\pi_{\theta},r\rangle\right\|=\left\|\nabla\ell(\theta)\right\|$
Hence, the loss for the bandit problem satisfies Assumption 5 with $L_{c}=0$ and $L_{g}=3$. Using Lemma 5 with $n=1$, we can conclude the the loss for the bandit problem satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\frac{3\sqrt{2}}{\Delta}$. From Mei et al. (2020, Lemma 3), we know that,
$\left\|\nabla\ell(\theta)\right\|=\left\|\nabla_{\theta}\langle\pi_{\theta},r\rangle\right\|\geq\pi_{\theta}(a^{*})\,[r(a^{*})-\langle{\pi_{t}},r\rangle]=\pi_{\theta}(a^{*})\,f(\theta)$
Hence, the loss for the bandit problem satisfies Assumption 4 with $\mu(\theta)=\pi_{\theta}(a^{*})$. ∎
Proposition 6.
Consider an infinite-horizon discounted Markov decision process (MDP) defined by $\langle{\mathcal{S}},{\mathcal{A}},\mathcal{P},r,\rho,\gamma\rangle$, where ${\mathcal{S}}$ and ${\mathcal{A}}$ represent the states and actions, $\mathcal{P}:{\mathcal{S}}\times{\mathcal{A}}\rightarrow\Delta_{{\mathcal{S}}}$ is the transition probability function, $r:{\mathcal{S}}\times{\mathcal{A}}\rightarrow[0,1]$ is the reward function, $\rho\in\Delta_{{\mathcal{S}}}$ is the initial state distribution, and $\gamma\in[0,1)$ represents the discount factor. If $V^{\pi}(s):=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s]$ where $s_{t}\sim p(.|s_{t-1},a_{t-1})$, and $a_{t}\sim\pi(.|s_{t})$ for $t\geq 1$ is the expected discounted cumulative reward for a policy $\pi$ starting at state $s$, we define $V^{\pi}(\rho):=\mathbb{E}_{s\sim\rho}[V^{\pi}(s)]$.
Consider a policy $\pi_{\theta}$ parameterized by $\theta\in\mathbb{R}^{|{\mathcal{S}}|\times|{\mathcal{A}}|}$ s.t. $\pi_{\theta}(s,\cdot)\in\Delta_{K}$ for all $s\in{\mathcal{S}}$ and $\pi_{\theta}(s,a)\propto\exp(\theta(s,a))$. The loss corresponding to the tabular MDP problem is given by:
$f(\theta)=V^{\pi^{*}}(\rho)-V^{\pi_{\theta}}(\rho)\,,$
where $\pi^{*}$ is the optimal policy. $f(\theta)$ satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\,\frac{\sqrt{2}}{\Delta\,(1-\gamma)}$,  Assumption 3 with $\nu=\frac{\sqrt{2}}{\Delta\,(1-\gamma)}$ and $\omega=0$ and  Assumption 4 with $\mu(\theta)=\frac{\min_{s}\pi_{\theta}(\pi^{*}(s)|s)}{\sqrt{S}\,\min_{s\in{\mathcal{S}}}\rho(s)}$.
Proof.
From Mei et al. (2020, Lemma 28), we know that if $\Delta:=\min_{s\in{\mathcal{S}}}Q^{*}(s,\pi^{*}(s))-\max_{a\neq\pi^{*}(s)}Q(s,a)$, then,
$\left\|\nabla f(\theta)\right\|\leq\frac{\sqrt{2}}{\Delta\,(1-\gamma)}\,f(\theta)$
Hence, the loss for the tabular MDP problem satisfies Assumption 3 with $\nu=\frac{\sqrt{2}}{\Delta\,(1-\gamma)}$ and $\omega=0$. Assuming that the starting state distribution has full support, i.e. $\rho(s)>0$, from Mei et al. (2020, Lemma 6), we know that,
$\left\|\nabla^{2}f(\theta)\right\|$ $\leq\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\cdot\sqrt{S}\cdot\left\|\nabla f(\theta)\right\|$
Hence, the loss for the tabular MDP problem satisfies Assumption 5 with $L_{c}=0$ and $L_{g}=\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]$. Using Lemma 5 with $n=1$, we can conclude the the loss for the tabular MDP problem satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\,\frac{\sqrt{2}}{\Delta\,(1-\gamma)}$. From Mei et al. (2020, Lemma 8), we know that
$\left\|\nabla f(\theta)\right\|$ $\geq\frac{\min_{s}\pi_{\theta}(\pi^{*}(a)|s)}{\sqrt{S}\,\min_{s\in{\mathcal{S}}\rho(s)}}\,f(\theta)$
Hence, the loss for the tabular MDP problem satisfies Assumption 4 with $\mu(\theta)=\frac{\min_{s}\pi_{\theta}(\pi^{*}(s)|s)}{\sqrt{S}\,\min_{s\in{\mathcal{S}}}\rho(s)}$. ∎
Proposition 7.
Consider the logistic regression objective in Eq. 3 with $n=2$ and $d=1$. Consider the two points to be such that $y_{1}\,x_{1}=2$ and $y_{2}\,x_{2}=-2$. For this problem, the non-uniformness assumption in Zhang et al. (2019): $\left\|\nabla^{2}f(\theta)\right\|\leq L_{0}+L_{1}\,\left\|\nabla f(\theta)\right\|$ cannot hold for $L_{0}=0$ and any $L_{1}\neq 0$.
Proof.
Using the proof of Proposition 1 to calculate the gradient and hessian, we get that $\nabla f_{1}(0)=0$ and $\nabla f_{2}(0)=0$ which implies $\nabla f(0)=0$. Similarly, for Hessian, we get $\nabla^{2}f_{1}(0)=1$ and $\nabla^{2}f_{2}(0)=1$ which implies $\nabla^{2}f(0)=1$. Since $\nabla f(\theta)=0$ and $\nabla^{2}f(\theta)\neq 0$, the assumption cannot hold with $L_{0}\neq 0$ and any $L_{1}\neq 0$. ∎
Appendix B Proofs for Section 3
See 1
Proof.
Case 1: If $L_{1}=0$Assumption 2 is equivalent to the standard $L_{0}$-uniform smoothness condition. In this case, we can follow the standard analysis of GD-LS (Nocedal & Wright, 2006) and conclude that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{2\,(1-c)}{L_{0}}\right\}\geq\min\left\{\eta_{\max},\frac{(1-c)}{3\,L_{0}}\right\}$.
In this special case, $\lambda_{0}=\frac{3\,L_{0}}{1-c}$ and $\lambda_{1}=0$, meaning that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}\right\}$. This concludes the proof.
Case 2: If $L_{1}\neq 0$ and since $f(\theta)$ is non-negative, we define the log-loss as follows.
$g(\theta):=\ln(L_{0}+L_{1}\,f(\theta))$
Using Assumption 2, $\nabla^{2}f(\theta)\preceq[L_{0}+L_{1}\,f(\theta)]\,I_{d}$. Using this result, we bound the Hessian of $g(\theta)$.
$\nabla g(\theta)$ $=\frac{L_{1}\,\nabla f(\theta)}{L_{0}+L_{1}\,f(\theta)}$
$\nabla^{2}g(\theta)$ $=\frac{L_{1}\,\nabla^{2}f(\theta)}{L_{0}+L_{1}\,f(\theta)}-\frac{L_{1}^{2}\,[\nabla f(\theta)][\nabla f(\theta)]^{T}}{(L_{0}+L_{1}\,f(\theta))^{2}}\preceq\frac{L_{1}\,\nabla^{2}f(\theta)}{L_{0}+L_{1}\,f(\theta)}$ (Since the second term is PSD)
$\implies\nabla^{2}g(\theta)\preceq L_{1}\,I_{d}$
Hence, $g(\theta)$ is $L_{1}$-globally smooth. Using this result, we know that for all $u,v$,
$g(u)$ $\leq g(v)+\langle\nabla g(v),u-v\rangle+\frac{L_{1}}{2}\,\left\|u-v\right\|_{2}^{2}$
Using this result for $u=\theta_{t+1}$ and $v=\theta_{t}$,
$g(\theta_{t+1})$ $\leq g(\theta_{t})+\langle\nabla g(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{L_{1}}{2}\,\left\|\theta_{t+1}-\theta_{t}\right\|_{2}^{2}$
$=g(\theta_{t})-{\eta_{t}}\,\langle\nabla f(\theta_{t}),\nabla g(\theta_{t})\rangle+\frac{L_{1}\,{\eta_{t}}^{2}}{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$
$=g(\theta_{t})-{\eta_{t}}\,\left\langle\nabla f(\theta_{t}),\frac{L_{1}\,\nabla f(\theta_{t})}{L_{0}+L_{1}\,f(\theta_{t})}\right\rangle+\frac{L_{1}\,{\eta_{t}}^{2}}{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$
$\implies g(\theta_{t}-{\eta_{t}}\,\nabla f(\theta_{t}))$ $\leq\underbrace{g(\theta_{t})-{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}+\frac{L_{1}\,{\eta_{t}}^{2}}{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}_{:=h_{Q}({\eta_{t}})}$ (10)
Next, we will compare the above inequality with what we obtain from the Armijo line-search.
$f(\theta_{t}-{\eta_{t}}\nabla f(\theta_{t}))$ $\leq f(\theta_{t})-c{\eta_{t}}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$
$L_{0}+L_{1}\,f(\theta_{t}-{\eta_{t}}\nabla f(\theta_{t}))$ $\leq L_{0}+L_{1}\,f(\theta_{t})-c{\eta_{t}}\,L_{1}\ \left\|\nabla f(\theta_{t})\right\|_{2}^{2}$
$\implies\ln\left(L_{0}+L_{1}\,f(\theta_{t}-{\eta_{t}}\nabla f(\theta_{t}))\right)$ $\leq\ln\left(L_{0}+L_{1}\,f(\theta_{t})-c{\eta_{t}}\,L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\right)$
$\implies g(\theta_{t}-{\eta_{t}}\nabla f(\theta_{t}))$ $\leq\ln\left(L_{0}+L_{1}\,f(\theta_{t})-c{\eta_{t}}\,L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\right)$
$=\ln\left((L_{0}+L_{1}\,f(\theta_{t}))\,\left(1-c{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}\right)\right)=g(\theta_{t})+\ln\left(1-c{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}\right)$
$\leq g(\theta_{t})+\left(1-c{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}\right)-1$
$\implies g(\theta_{t+1})$ $\leq\underbrace{g(\theta_{t})-c{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}}_{:=h_{L}({\eta_{t}})}$ (11)
Hence, assuming exact back-tracking, if ${\eta_{t}}$ is a step-size that satisfies Eq. 6, then Eq. 11 will also be satisfied.
If the Armijo condition is satisfied for an ${\eta_{t}}$ s.t. $h_{L}({\eta_{t}})\leq h_{Q}({\eta_{t}})$, then,
$g(\theta_{t})-c{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}$ $\leq g(\theta_{t})-{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}+\frac{L_{1}\,{\eta_{t}}^{2}}{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$
$\implies{\eta_{t}}\geq\frac{2\,(1-c)}{L_{0}+L_{1}\,f(\theta_{t})}$
If the Armijo condition is satisfied for an ${\eta_{t}}$ s.t. $h_{Q}({\eta_{t}})\leq h_{L}({\eta_{t}})$, it implies that ${\eta_{t}}\leq\frac{2(1-c)}{L_{0}+L_{1}f(\theta_{t})}$. However, we show that the resulting step-size cannot be too small. In particular, we will prove that the Armijo condition is satisfied for ${\eta_{t}}=\frac{2(1-c)}{6(L_{0}+L_{1}\omega)+6\,L_{1}\,(\nu+1)\,f(\theta_{t})}$. To show this, we use Assumption 2. In order to use this inequality, we have to ensure that $\left\|\theta_{t+1}-\theta_{t}\right\|={\eta_{t}}\,\left\|\nabla f(\theta_{t})\right\|\leq\frac{q}{L_{1}}$. Since based on Assumption 3, $\left\|\nabla f(\theta_{t})\right\|\leq\nu\,f(\theta_{t})+\omega$, it suffices to ensure that $q\geq{\eta_{t}}\,L_{1}(\nu\,f(\theta_{t})+\omega)$.
Using Lemma 5, we get that:
$f(\theta_{t+1})\leq f(\theta_{t})-{\eta_{t}}\left\|\nabla f(\theta_{t})\right\|_{2}^{2}+\frac{AL_{0}+B\,L_{1}\,f(\theta_{t})}{2}\,{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$
The Armijo condition is definitely satisfied if:
$f(\theta_{t})-{\eta_{t}}\left\|\nabla f(\theta_{t})\right\|_{2}^{2}+\frac{\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})}{2}\,{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\leq f(\theta_{t})-c{\eta_{t}}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$
Hence, the Armijo condition is satisfies for all ${\eta_{t}}$ s.t.
$\implies{\eta_{t}}$ $\leq\frac{2\,(1-c)}{\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})}$
Since,
$\frac{2\,(1-c)}{\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,(\nu f(\theta_{t})+\omega)}\leq\frac{2\,(1-c)}{\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})}\,,$
the Armijo condition will be satisfied for the smaller step-size.
Moreover, for ${\eta_{t}}^{\prime}:=\frac{2(1-c)}{\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,(\nu f(\theta_{t})+\omega)}$, we need to ensure that $q\geq{\eta_{t}}^{\prime}\,L_{1}(\nu\,f(\theta_{t})+\omega)$. Hence, we want to find a $q$ s.t.
$q$ $\geq\frac{2(1-c)\,L_{1}(\nu\,f(\theta_{t})+\omega)}{\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,(\nu f(\theta_{t})+\omega)}$
Since $\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})>0$, it suffices to choose $q$ s.t.
$\implies q$ $\geq\frac{2(1-c)\,L_{1}(\nu\,f(\theta_{t})+\omega)}{\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,(\nu f(\theta_{t})+\omega)}=\frac{2(1-c)}{\left(\frac{e^{q}-1}{q}\right)}$
Finally, since $1+x\leq\exp(x)$ for all $x$, it suffices to choose $q$ s.t.
$q$ $\geq 2(1-c)$ (since $x+1\leq\exp(x)$)
Hence, $q=2$ satisfies the required conditions. Therefore, for $q=2$ we have,
${\eta_{t}}^{\prime}=\frac{2(1-c)}{\left(1+e^{2}-\frac{e^{2}-1}{2}\right)\,L_{0}+\left(\frac{e^{2}-1}{2}\right)\,L_{1}\,f(\theta_{t})+\left(\frac{e^{2}-1}{2}\right)\,L_{1}\,(\nu f(\theta_{t})+\omega)}$
Therefore for any ${\eta_{t}}\leq{\eta_{t}}^{\prime}$, we have $q=2\geq{\eta_{t}}L_{1}\,(\nu f(\theta_{t})+\omega)$. Since $\frac{e^{2}-1}{2}\leq 6$ and $1+e^{2}-\frac{e^{2}-1}{2}\leq 6$ we can set
${\eta_{t}}=\frac{2(1-c)}{6\,L_{0}+6\,L_{1}\,f(\theta_{t})+6\,L_{1}\,(\nu f(\theta_{t})+\omega)}=\frac{2(1-c)}{6\,(L_{0}+L_{1}\omega)+6\,L_{1}(\nu+1)\,f(\theta_{t})}.$
Based on above argument, we can conclude that the ${\eta_{t}}$, the step-size returned by the Armijo line-search is lower-bounded as ${\eta_{t}}\geq\frac{2(1-c)}{6\,(L_{0}+L_{1}\omega)+6\,L_{1}(\nu+1)\,f(\theta_{t})}$. Moreover if $\eta_{\max}\leq\frac{2(1-c)}{6\,(L_{0}+L_{1}\omega)+6\,L_{1}(\nu+1)\,f(\theta_{t})}$, then $\eta_{\max}$ satisfies both Armijo condition and $h_{Q}(\eta_{\max})\leq h_{L}(\eta_{\max})$, in which case, the line-search would terminate immediately and return $\eta_{\max}$. Therefore ${\eta_{t}}\geq\min\{\eta_{\max},\frac{2(1-c)}{6\,(L_{0}+L_{1}\omega)+6\,L_{1}(\nu+1)\,f(\theta_{t})}\}$. ∎
See 1
Proof.
Using the Armijo line-search condition in Eq. 6, and combining it with the lower-bound in Lemma 1,
$f(\theta_{t+1})$ $\leq f(\theta_{t})-\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ (12)
We now follow a proof similar to that of Axiotis & Sviridenko (2023, Theorem 5.2) and derive a linear rate of convergence. From the theorem assumption, we know that $\left\|\nabla f(\theta)\right\|_{2}^{2}\geq\frac{[f(\theta)-f^{*}]^{2}}{R}$. Combining these relations,
$f(\theta_{t+1})$ $\leq f(\theta_{t})-\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}\,\frac{[f(\theta_{t})-{f^{*}}]^{2}}{R}$
Let us define $\tau:=\max\{t\,\text{ s.t }\,\lambda_{0}\leq\lambda_{1}f(\theta_{t})\}$. Hence, for all $t\leq\tau$, $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$.
Consider two cases:
Case (1): If ${f^{*}}\geq\frac{\lambda_{0}}{\lambda_{1}}-\epsilon$. Since $\{f(\theta_{t})\}_{t=0}^{t=\tau}$ is monotonically decreasing due to the Armijo line-search and converging to $\frac{\lambda_{0}}{\lambda_{1}}$. Hence, there exists a $\tau^{\prime}$ s.t. $\tau^{\prime}\leq\tau$ such that $f(\theta_{\tau^{\prime}})-{f^{*}}\leq\epsilon$ and $f(\theta_{\tau^{\prime}-1})-{f^{*}}\geq\epsilon$, i.e. $\tau^{\prime}$ is the iteration index when the desired sub-optimality criterion is satisfied for the first time. This implies that for all $t<\tau^{\prime}$, $\delta_{t}:=f(\theta_{t})-{f^{*}}>\epsilon$. Hence, $\frac{f(\theta_{\tau^{\prime}})}{{f^{*}}}\leq\gamma:=1+\frac{\epsilon}{{f^{*}}}$. Since ${f^{*}}>0$ and $\epsilon>0$, $\gamma>1$. Hence for all $t<\tau^{\prime}$,
$\frac{f(\theta_{t})}{{f^{*}}}>\gamma\implies\frac{\delta_{t}}{f(\theta_{t})}=1-\frac{{f^{*}}}{f(\theta_{t})}>1-\frac{1}{\gamma}>0.$
Using the condition of Case (1), we get
$\delta_{t+1}$ $\leq\delta_{t}-\underbrace{\frac{1}{2\lambda_{1}\,R}}_{:=\alpha}\frac{[f(\theta_{t})-{f^{*}}]^{2}}{f(\theta_{t})}$
$\leq\delta_{t}-\bar{\alpha}\frac{[f(\theta_{t})-{f^{*}}]^{2}}{f(\theta_{t})}$ (where $\bar{\alpha}:=\max\{1,\alpha\}$)
$\leq\delta_{t}-\bar{\alpha}\frac{[f(\theta_{t})-{f^{*}}]}{f(\theta_{t})}\delta_{t}$
$=\delta_{t}-\bar{\alpha}\left(1-\frac{{f^{*}}}{f(\theta_{t})}\right)\delta_{t}$
Combining the above relations, for all $t<\tau^{\prime}$,
$\delta_{t+1}$ $\leq\left(1-\underbrace{\bar{\alpha}\,\left(1-\frac{1}{\gamma}\right)}_{:=\rho}\right)\delta_{t}$
Since $\bar{\alpha}\in(0,1)$ and $\left(1-\frac{1}{\gamma}\right)\in(0,1)$, $\rho:=\bar{\alpha}\,\left(1-\frac{1}{\gamma}\right)\in(0,1)$. Recursing from $t=0$ to $t=\tau^{\prime}-1$,
$\delta_{\tau^{\prime}}$ $\leq\exp\left(-\rho\,\tau^{\prime}\right)\,\delta_{0}$
In order to ensure that $f(\theta_{\tau^{\prime}})-{f^{*}}\leq\epsilon$, we require,
$\tau^{\prime}\geq\frac{1}{\rho}\ln\left(\frac{\delta_{0}}{\epsilon}\right)=\frac{1}{\min\{\alpha,1\}}\,\left(\frac{{f^{*}}}{\epsilon}+1\right)\,\ln\left(\frac{\delta_{0}}{\epsilon}\right)=\max\{2\,R\lambda_{1},1\}\,\left(\frac{{f^{*}}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-{f^{*}}}{\epsilon}\right)$
Case (2): If ${f^{*}}<\frac{\lambda_{0}}{\lambda_{1}}-\epsilon$. We will divide the subsequent analysis into two phases.
Phase (1): For all $t\leq\tau$, s.t. $\lambda_{0}+\lambda_{1}\,f(\theta_{t})\leq 2\lambda_{1}\,f(\theta_{t})$ holds, by a similar analysis as above, we can conclude that,
$\delta_{\tau}$ $\leq\exp\left(-\rho\,\tau\right)\,\delta_{0}\implies f(\theta_{\tau})-{f^{*}}\leq\exp\left(-\rho\,\tau\right)\,[f(\theta_{0})-{f^{*}}]$
Since $\delta_{\tau}=f(\theta_{\tau})-{f^{*}}\geq\frac{\lambda_{0}}{\lambda_{1}}-{f^{*}}=\epsilon$. Hence,
$\exp\left(-\rho\,\tau\right)\,[f(\theta_{0})-{f^{*}}]\geq\epsilon\implies\tau\leq\frac{1}{\rho}\,\ln\left(\frac{f(\theta_{0})-{f^{*}}}{\epsilon}\right)$
Phase (2): For all $t>\tau$, $\lambda_{0}\geq\lambda_{1}\,f(\theta_{t})$ which implies $\lambda_{0}+\lambda_{1}\,f(\theta_{t})\leq 2\lambda_{0}$. In this case,
$f(\theta_{t+1})-{f^{*}}$ $\leq\underbrace{[f(\theta_{t})-{f^{*}}]}_{:=\delta_{t}}-\frac{1}{2\lambda_{0}R}\,[f(\theta_{t})-{f^{*}}]^{2}\implies\delta_{t+1}\leq\delta_{t}-\frac{1}{2\lambda_{0}\,R}\delta_{t}^{2}$
Following the standard approach, we divide both sides by $\delta_{t+1}\delta_{t}$ and rearranging we get
$\frac{1}{2\lambda_{0}\,R}$ $\leq\frac{1}{2\lambda_{0}\,R}\frac{\delta_{t}}{\delta_{t+1}}$ (since $\frac{\delta_{t}}{\delta_{t+1}}\geq 1$)
$\leq\frac{1}{\delta_{t+1}}-\frac{1}{\delta_{t}}$
Summing the above for $t=\tau$ to $t=T-1$, we get
$\frac{T-\tau}{2\lambda_{0}\,R}\leq\frac{1}{\delta_{T}}-\frac{1}{\delta_{\tau}}$
$\implies\delta_{T}\leq\frac{1}{\frac{T-\tau}{2\lambda_{0}\,R}+\frac{1}{\delta_{\tau}}}$
We need to find $T$ such that $\delta_{T}\leq\epsilon$, which means
$\frac{1}{\frac{T-\tau}{2\lambda_{0}\,R}+\frac{1}{\delta_{\tau}}}\leq\epsilon\implies T-\tau$ $\geq\frac{2\lambda_{0}\,R}{\epsilon}-\frac{2\lambda_{0}\,R}{\delta_{\tau}}\implies T\geq\frac{2\lambda_{0}\,R}{\epsilon}+\frac{1}{\rho}\ln(\delta_{0}/\epsilon)$
Putting everything together,
$T$ $\geq\frac{2\lambda_{0}\,R}{\epsilon}+\max\{2\,R\lambda_{1},1\}\,\left(\frac{{f^{*}}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-{f^{*}}}{\epsilon}\right)$
Appendix C Proofs for Section 4
See 1
Proof.
Using the convexity of $f$,
$f(\theta_{t})-f^{*}$ $\leq\langle\nabla f(\theta_{t}),\theta_{t}-\theta^{*}\rangle\leq\left\|\nabla f(\theta_{t})\right\|\,\left\|\theta_{t}-\theta^{*}\right\|$
$\implies\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ $\geq\frac{[f(\theta_{t})-f^{*}]^{2}}{\left\|\theta_{t}-\theta^{*}\right\|_{2}^{2}}$
Next, we show that $\left\|\theta_{t+1}-\theta^{*}\right\|\leq\left\|\theta_{t}-\theta^{*}\right\|$ for all $t$, and hence $\left\|\theta_{t}-\theta^{*}\right\|\leq\left\|\theta_{0}-\theta^{*}\right\|$.
$\left\|\theta_{t+1}-\theta^{*}\right\|_{2}^{2}$ $=\left\|\theta_{t}-\theta^{*}-{\eta_{t}}\nabla f(\theta_{t})\right\|_{2}^{2}=\left\|\theta_{t}-\theta^{*}\right\|_{2}^{2}-2{\eta_{t}}\langle\nabla f(\theta_{t}),\theta_{t}-\theta^{*}\rangle+{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$
$\leq\left\|\theta_{t}-\theta^{*}\right\|_{2}^{2}-2{\eta_{t}}[f(\theta_{t})-f^{*}]+{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ (By convexity of $f$)
$\leq\left\|\theta_{t}-\theta^{*}\right\|_{2}^{2}-2{\eta_{t}}[f(\theta_{t})-f^{*}]+2\,{\eta_{t}}\,[f(\theta_{t})-f(\theta_{t+1})]$ (Using the Armijo line-search with $c=\frac{1}{2}$)
$\implies\left\|\theta_{t+1}-\theta^{*}\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-\theta^{*}\right\|_{2}^{2}-2{\eta_{t}}\,[f(\theta_{t+1})-f^{*}]+2\,{\eta_{t}}\,[f(\theta_{t})-{f^{*}}]$
$\leq\left\|\theta_{t}-\theta^{*}\right\|_{2}^{2}$ (By the definition of $\theta^{*}$ and using that for all $t$, $f^{*}<f(\theta_{T})\leq f(\theta_{t})$)
Combining the above inequalities,
$\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ $\geq\frac{[f(\theta_{t})-f^{*}]^{2}}{\left\|\theta_{0}-\theta^{*}\right\|_{2}^{2}}$ (13)
Using Theorem 1 with $R=\left\|\theta_{0}-\theta^{*}\right\|_{2}^{2}$ and setting $L_{0}=0$ completes the proof. ∎
See 2
Proof.
For an arbitrary comparator $u$ s.t. $f(u)\leq(\theta_{T})$,
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $=\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,\langle\nabla f(\theta_{t}),\theta_{t}-u\rangle+{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ (Convexity)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+\frac{{\eta_{t}}}{c}\,[f(\theta_{t})-f(\theta_{t+1})]$ (Using the Armijo line-search with $c>\frac{1}{2}$)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+\frac{{\eta_{t}}}{c}\,[f(\theta_{t})-f(u)]$ (Since $f(u)\leq f(\theta_{t})$)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-(2-\frac{1}{c})\,{\eta_{t}}\,[f(\theta_{t})-f(u)]$ (14)
$\implies\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-(2-\frac{1}{c})\,\frac{1}{\lambda_{0}+\lambda_{1}f(\theta_{t})}[f(\theta_{t})-f(u)]$ (Using Lemma 1)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-(2-\frac{1}{c})\,\frac{1}{\lambda_{1}}\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}$ ( using $\lambda_{0}=0$ since $L_{0},\omega=0$ and by defining $C:=(2-\frac{1}{c})\,(1/\lambda_{1})$)
By recursing from $t=0$ to $T-1$,
$\left\|\theta_{T}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-C\,\sum_{t=0}^{T-1}\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}$
Assume $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon$. Hence, $\frac{f(\theta_{T})}{f(u)}\leq\gamma:=1+\frac{\epsilon}{f(u)}$. Hence, for all $t<T$, $f(\theta_{t})-f(u)>\epsilon$ and $\frac{f(\theta_{t})}{f(u)}>\gamma$. Consequently, $\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\geq 1-\frac{1}{\gamma}$. Combining the above relations,
$\left\|\theta_{T}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\,\left(1-\frac{1}{\gamma}\right)$
Since $f$ is $L$-uniform smooth, we know that $f(\theta_{T})-f(u)\leq\frac{L}{2}\left\|\theta_{T}-u\right\|_{2}^{2}$. Hence, for $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to guarantee that $\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{2\,\epsilon}{L}$. In order to guarantee this, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{2\epsilon}{L}}{C}\,\left[1+\frac{f(u)}{\epsilon}\right]$ (Using the definition of $\gamma$)
Using the definition of $C$, we conclude that it is sufficient to set $T$ as:
$T$ $\geq\frac{c\lambda_{1}\,\left\|\theta_{0}-u\right\|_{2}^{2}}{(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right]$
See 2
Proof.
Define $u^{*}$ to be the max-margin solution i.e. $\left\|u^{*}\right\|=1$ and $\gamma$ to be the corresponding margin, i.e.
$\gamma$ $:=\min_{i}\,y_{i}\langle x_{i},u^{*}\rangle$ (15)
For a scalar $\beta>0$,
$f(\beta u^{*})$ $=\frac{1}{n}\,\sum_{i=1}^{n}\ln(1+\exp(-y_{i}\langle x_{i},\beta u^{*}\rangle))\leq\frac{1}{n}\,\sum_{i=1}^{n}\exp(-y_{i}\langle x_{i},u^{*}\rangle)\leq\exp(-\beta\gamma)$ (16)
For normalized data, s.t. $\left\|x_{i}\right\|\leq 1$, the logistic regression loss is convex, is uniform smooth with $L=\frac{\lambda_{\max}[X^{T}X]}{4n}\leq 1$.
We set $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$ implies that $f(\beta u^{*})\leq\epsilon$. For the $T$ defined in the theorem statement, consider two cases:
Case (I): $f(\theta_{T})<f(\beta u^{*})\leq\epsilon$. This gives the desired result immediately.
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, we can use the result in Theorem 2. In particular, for a comparator $u=\beta u^{*}$, GD with Armijo line-search with $c$, $\eta_{\max}=\infty$ and $\theta_{0}=0$ ensures that when $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon\implies f(\theta_{T})\leq 2\epsilon$, then, for $C:=\frac{2c-1}{c\lambda_{1}}$,
$f(\theta_{T})$ $\leq f(\beta u^{*})+\frac{L}{2}\left[\beta^{2}-C\,T\,\left(\frac{\epsilon}{\epsilon+f(\beta u^{*})}\right)\right]$
$f(\theta_{T})$ $\leq\epsilon+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}-\frac{C\,T}{2}\right]$
Hence, in order to ensure that $f(\theta_{T})\leq 2\,\epsilon$, it is sufficient to set $T$ as:
$T$ $\geq\frac{1}{C\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
$=\frac{c\lambda_{1}}{(2c-1)\,\gamma^{2}}\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
$=\frac{3\,c\,L_{1}\,(\nu+1)}{(2c-1)(1-c)\gamma^{2}}\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$ (using the value of $\lambda_{1}$)
$=\frac{6\,c\,}{(2c-1)(1-c)\gamma^{2}}\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$ (using Proposition 1 for the value of $\nu$ and $L_{1}$)
Combining the two cases, we conclude that $f(\theta_{T})\leq 2\epsilon$. ∎
C.1 Proofs for the Polyak Step-size
For an arbitrary comparator $u$, we generalize the Polyak step-size (Polyak, 1987) at iteration $t\in[T]$ as:
${\eta_{t}}$ $=\frac{f(\theta_{t})-f(u)}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}\,,$ (17)
where, $c\in(0,1)$ is a hyper-parameter. Note that when $u=\theta^{*}=\operatorname*{arg\,min}f(\theta)$, ${\eta_{t}}=\frac{f(\theta_{t})-f^{*}}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ recovers the standard Polyak step-size in Polyak (1987).
We analyze the convergence of GD with the step-size in Eq. 17 under Assumption 3 with $\omega=0$ i.e. we will assume that $f$ is $L$ uniform smooth and that for all $\theta$, $\left\|\nabla f(\theta)\right\|\leq\nu\,f(\theta)$. From Propositions 1, 5 and 4, we know that this property is true from binary classification with the logistic loss, as well as for multi-class classification with the cross-entropy loss.
Theorem 4.
For a fixed $\epsilon>0$, assuming $f(\theta)$ is convex, $L$-uniform smooth and satisfies Assumption 3 with $\omega=0$, GD with the Polyak step-size in Eq. 17 and $c>\frac{1}{2}$, requires
$T$ $\geq\frac{c^{2}\,\nu\,\left\|\theta_{0}-u\right\|_{2}^{2}}{(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right]^{2}$
iterations to ensure that $f(\theta_{T})-f(u)\leq\epsilon$, where $u$ is an arbitrary comparator s.t. $f(u)<f(\theta_{T})$.
Proof.
Following a proof similar to that for Theorem 2, for an arbitrary comparator $u$ s.t. $f(u)\leq(\theta_{T})$,
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $=\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,\langle\nabla f(\theta_{t}),\theta_{t}-u\rangle+{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ (Convexity)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+\frac{{\eta_{t}}}{c}\,[f(\theta_{t})-f(u)]$ (Using the Polyak step-size in Eq. 17 with $c>\frac{1}{2}$ to simplify the third term)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-f(u)]^{2}}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ (Using the Polyak step-size in Eq. 17)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-f(u)]^{2}}{c\,\nu^{2}\,[f(\theta_{t})]^{2}}$ (Using Assumption 3)
$\implies\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right)^{2}$ (Using Lemma 1 and defining $C:=\frac{(2c-1)}{c^{2}\,\nu}$)
By recursing from $t=0$ to $T-1$,
$\left\|\theta_{T}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-C\,\sum_{t=0}^{T-1}\left(\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right)^{2}$
Assume $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon$. Hence, $\frac{f(\theta_{T})}{f(u)}\leq\gamma:=1+\frac{\epsilon}{f(u)}$. Hence, for all $t<T$, $f(\theta_{t})-f(u)>\epsilon$ and $\frac{f(\theta_{t})}{f(u)}>\gamma$. Consequently, $\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\geq 1-\frac{1}{\gamma}$. Combining the above relations,
$\left\|\theta_{T}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\,\left(1-\frac{1}{\gamma}\right)^{2}$
Since $f$ is $L$-uniform smooth, we know that $f(\theta_{T})-f(u)\leq\frac{L}{2}\left\|\theta_{T}-u\right\|_{2}^{2}$. Hence, for $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to guarantee that $\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{2\,\epsilon}{L}$. In order to guarantee this, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{2\epsilon}{L}}{C}\,\left[1+\frac{f(u)}{\epsilon}\right]^{2}$ (Using the definition of $\gamma$)
Using the definition of $C$, we conclude that it is sufficient to set $T$ as:
$T$ $\geq\frac{c^{2}\,\nu\,\left\|\theta_{0}-u\right\|_{2}^{2}}{(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right]^{2}$
We use the above result and prove the following corollary for logistic regression on separable data.
Corollary 6.
For logistic regression on linearly separable data with margin $\gamma$ where $u^{*}$ is the corresponding max-margin solution and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon>0$, GD with the Polyak step-size ${\eta_{t}}=\frac{f(\theta_{t})-f(\beta u^{*})}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ requires
$T\geq\frac{4\,c^{2}}{(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
to ensure that $f(\theta_{T})\leq 2\,\epsilon$.
Proof.
Define $u^{*}$ to be the max-margin solution i.e. $\left\|u^{*}\right\|=1$ and $\gamma$ to be the corresponding margin, i.e.
$\gamma$ $:=\min_{i}\,y_{i}\langle x_{i},u^{*}\rangle$ (18)
For a scalar $\beta>0$,
$f(\beta u^{*})$ $=\frac{1}{n}\,\sum_{i=1}^{n}\ln(1+\exp(-y_{i}\langle x_{i},\beta u^{*}\rangle))\leq\frac{1}{n}\,\sum_{i=1}^{n}\exp(-y_{i}\langle x_{i},u^{*}\rangle)\leq\exp(-\beta\gamma)$ (19)
For normalized data, s.t. $\left\|x_{i}\right\|\leq 1$, the logistic regression loss is convex, is uniform smooth with $L=\frac{\lambda_{\max}[X^{T}X]}{4n}\leq 1$. Moreover, $\nu=1$. We set $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$ implies that $f(\beta u^{*})\leq\epsilon$. For the $T$ defined in the theorem statement, consider two cases:
Case (I): $f(\theta_{T})<f(\beta u^{*})\leq\epsilon$. This gives the desired result immediately.
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, we can use the result in Theorem 4. In particular, for a comparator $u=\beta u^{*}$, GD with the Polyak step-size and $\theta_{0}=0$ ensures that when $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon\implies f(\theta_{T})\leq 2\epsilon$, then, for $C:=\frac{(2c-1)}{c^{2}\,\nu}=\frac{(2c-1)}{c^{2}}$,
$f(\theta_{T})$ $\leq f(\beta u^{*})+\frac{L}{2}\left[\beta^{2}-C\,T\,\left(\frac{\epsilon}{\epsilon+f(\beta u^{*})}\right)^{2}\right]$
$f(\theta_{T})$ $\leq\epsilon+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}-\frac{C\,T}{4}\right]$
Hence, in order to ensure that $f(\theta_{T})\leq 2\,\epsilon$, it is sufficient to set $T$ as:
$T\geq\frac{4\,c^{2}}{(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
Combining the two cases, we conclude that $f(\theta_{T})\leq 2\epsilon$. ∎
In the above result, $\eta_{t}$ depends on unknown problem-dependent constants such as $u^{*}$ and $\gamma$. In the following corollary, we show that we can remove this dependence, at the expense of obtaining a worse iteration complexity.
Corollary 7.
For logistic regression on linearly separable data with margin $\gamma$ where $u^{*}$ is the corresponding max-margin solution and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$, if $\alpha:=\frac{f(\beta u^{*})}{\epsilon}$ and if for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon>0$, GD with the Polyak step-size ${\eta_{t}}=\frac{f(\theta_{t})-\epsilon}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ requires
$T\geq\frac{(1+\alpha)^{2}\,c^{2}}{\alpha^{2}\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
to ensure that $f(\theta_{T})\leq 2\,\epsilon$.
Proof.
The first part of the proof considers a general convex $f$ which is $L$ uniform-smooth and satisfies Assumption 3 with $\omega=0$. Assume $\alpha\in(0,1)$ and consider an arbitrary comparator $u$ s.t. $f(u)\leq f(\theta_{T})$ and $\alpha\epsilon\leq f(u)\leq\epsilon$. We set $\eta_{t}=\frac{f(\theta_{t})-\epsilon}{c\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$.
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $=\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,\langle\nabla f(\theta_{t}),\theta_{t}-u\rangle+{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ (20)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ (Convexity)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+\frac{{\eta_{t}}}{c}\,[f(\theta_{t})-\epsilon]$ (Using the Polyak step-size with $c>\frac{1}{2}$ to simplify the third term)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+\frac{{\eta_{t}}}{c}\,[f(\theta_{t})-f(u)]$ (Since $f(u)\leq\epsilon$)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\eta_{t}\,[f(\theta_{t})-f(u)]$ (21)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-\epsilon][f(\theta_{t})-f(u)]}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ (Using the Polyak step-size )
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-\epsilon]^{2}}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ (Since $f(\theta_{t})>\epsilon$ for all $t$. Else, we have achieved the desired sub-optimality)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-\epsilon]^{2}}{c\,\nu^{2}\,[f(\theta_{t})]^{2}}$ (Using Assumption 3)
$\implies\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(\frac{f(\theta_{t})-\epsilon}{f(\theta_{t})}\right)^{2}$ (Using Lemma 1 and defining $C:=\frac{(2c-1)}{c^{2}\,\nu}$)
Assuming $T$ is the first iterate s.t. $f(\theta_{T})-f(u)\leq\epsilon$, and therefore $f(\theta_{t})-f(u)\geq\epsilon$ that implies $f(\theta_{t})\geq\epsilon+f(u)\geq(1+\alpha)\epsilon$. Therefore
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(\frac{f(\theta_{t})-\epsilon}{f(\theta_{t})}\right)^{2}$
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(1-\frac{\epsilon}{(1+\alpha)\epsilon}\right)^{2}$
$=\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(1-\frac{1}{(1+\alpha)}\right)^{2}$
By recursing from $t=0$ to $T-1$,
$\left\|\theta_{T}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-CT\left(1-\frac{1}{(1+\alpha)}\right)^{2}$
Since $f$ is $L$-uniform smooth, we know that $f(\theta_{T})-f(u)\leq\frac{L}{2}\left\|\theta_{T}-u\right\|_{2}^{2}$. Hence, for $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to guarantee that $\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{2\,\epsilon}{L}$. In order to guarantee this, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{2\epsilon}{L}}{C}\,\left[1+\frac{1}{\alpha}\right]^{2}$
Using the definition of $C$, we conclude that it is sufficient to set $T$ as:
$T$ $\geq\frac{c^{2}\,\nu\,\left\|\theta_{0}-u\right\|_{2}^{2}}{(2c-1)}\,\left[1+\frac{1}{\alpha}\right]^{2}$ (22)
This completes the first part of the proof. The subsequent proof is specialized to unregularized logistic regression.
Define $u^{*}$ to be the max-margin solution i.e. $\left\|u^{*}\right\|=1$ and $\gamma$ to be the corresponding margin, i.e.
$\gamma$ $:=\min_{i}\,y_{i}\langle x_{i},u^{*}\rangle$ (23)
For a scalar $\beta>0$,
$f(\beta u^{*})$ $=\frac{1}{n}\,\sum_{i=1}^{n}\ln(1+\exp(-y_{i}\langle x_{i},\beta u^{*}\rangle))\leq\frac{1}{n}\,\sum_{i=1}^{n}\exp(-y_{i}\langle x_{i},u^{*}\rangle)\leq\exp(-\beta\gamma)$ (24)
For normalized data, s.t. $\left\|x_{i}\right\|\leq 1$, the logistic regression loss is convex, is uniform smooth with $L=\frac{1}{4n}\lambda_{\max}[X^{T}X]\leq 1$. Moreover, $\nu=1$. We set $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$ implies that $f(\beta u^{*})\leq\epsilon$. For the $T$ defined in the theorem statement, consider two cases:
Case (I): $f(\theta_{T})<f(\beta u^{*})\leq\epsilon$. This gives the desired result immediately.
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, $\alpha=\frac{f(\beta u^{*})}{\epsilon}\leq 1$ and we can use the above result. In particular, we choose $u=\beta u^{*}$ and $\theta_{0}=0$ ensuring that $\left\|\theta_{0}-u\right\|=\beta$. Plugging this value in Eq. 22, we conclude that in order to ensure that $f(\theta_{T})\leq 2\,\epsilon$, it is sufficient to set $T$ as:
$T\geq\frac{(1+\alpha)^{2}\,c^{2}}{\alpha^{2}\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
Combining the two cases, we conclude that $f(\theta_{T})\leq 2\epsilon$. ∎
Though the above result does not require the knowledge of any problem-dependent constants, the iteration complexity scales as $O\left(\frac{1}{\alpha^{2}}\,\ln\left(\frac{1}{\epsilon}\right)^{2}\right)$, where $\alpha=\frac{f(\beta\,u^{*})}{\epsilon}$ can be arbitrarily small though non-zero.
Appendix D Proofs for Section 5
D.1 Proofs for Section 5.1
See 3
Proof.
Using Assumption 4, we know that,
$\left\|\nabla f(\theta)\right\|_{2}^{2}\geq[\mu(\theta)]^{2}[f(\theta)-f^{*}]^{2}\geq\mu^{2}\,[f(\theta)-f^{*}]^{2}$
Using Theorem 1 with $R=\frac{1}{\mu^{2}}$ and $L_{0}=0$ completes the proof. ∎
D.2 Proofs for Section 5.2
See 3
Proof.
Using the Armijo line-search condition in Eq. 6, and combining it with the lower-bound in Lemma 1,
$f(\theta_{t+1})$ $\leq f(\theta_{t})-\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ (25)
Phase 1: Let us define $\tau:=\max\{t\,\text{ s.t }\,\lambda_{0}\leq\lambda_{1}\,f(\theta_{t})\}$. Hence, for all $t\leq\tau$, $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$, and hence,
$f(\theta_{t+1})$ $\leq f(\theta_{t})-\frac{1}{2\,\lambda_{1}\,f(\theta_{t})}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\,.$
From Assumption 4 with $\zeta=2$, we know that $\left\|\nabla f(\theta)\right\|_{2}^{2}\geq\mu(\theta)\,[f(\theta)-f^{*}]$ and that $f^{*}=0$. Combining these relations, we have that for all $t\leq\tau$,
$f(\theta_{t+1})$ $\leq f(\theta_{t})-\frac{f(\theta_{t})}{2\,\lambda_{1}\,R\,f(\theta_{t})}=f(\theta_{t})-\underbrace{\frac{\mu}{2\,\lambda_{1}}}_{:=\alpha}$ (Since $\mu=\min_{t\in[T]}\mu(\theta_{t})$)
Recursing from $t=0$ to $t=\tau$,
$f(\theta_{\tau})$ $\leq f(\theta_{0})-\tau\,\alpha$
Since $f(\theta_{\tau})\geq\frac{\lambda_{0}}{\lambda_{1}}$, we get that,
$\tau\leq\frac{1}{\alpha}\,\left[f(\theta_{0})-\frac{\lambda_{0}}{\lambda_{1}}\right]=\frac{2\lambda_{1}}{\mu}\,\left[f(\theta_{0})-\frac{\lambda_{0}}{\lambda_{1}}\right]$
Phase 2: For $t\geq\tau$, $f(\theta_{t})\leq\frac{\lambda_{0}}{\lambda_{1}}$. Combining this with Eq. 25 and using that $\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\geq\mu\,f(\theta_{t})$,
$f(\theta_{t+1})$ $\leq f(\theta_{t})-\frac{\mu\,f(\theta_{t})}{2\,\lambda_{0}}=f(\theta_{t})\left(1-\frac{\mu}{2\,\lambda_{0}}\right)$
Recursing from $t=\tau$ to $t=T$ and using that $1-x\leq\exp(-x)$,
$f(\theta_{T})$ $\leq\exp\left(-\frac{\mu\,(T-\tau)}{2\,\lambda_{0}}\right)\,f(\theta_{\tau})$
Hence, in order for $f(\theta_{T})\leq\epsilon$, we require
$T$ $\geq\tau+2\,\frac{\lambda_{0}}{\mu}\,\ln\left(\frac{f(\theta_{\tau})}{\epsilon}\right)$
Since $f(\theta_{\tau})\leq f(\theta_{0})-\tau\,\alpha$, is sufficient to set $T$ as:
$T$ $\geq\underbrace{\tau+2\,\frac{\lambda_{0}}{\mu}\,\ln\left(\frac{f(\theta_{0})-\tau\,\alpha}{\epsilon}\right)}_{:=h(\tau)}$
Hence, it is sufficient to set $T$ as:
$T$ $\geq\max_{\tau}h(\tau)\quad\text{s.t}\quad\tau\leq 2\frac{\lambda_{1}}{\mu}\,\,\left[f(\theta_{0})-\frac{\lambda_{0}}{\lambda_{1}}\right]\,.$
Calculating the first and second derivatives of $h(\tau)$,
$h^{\prime}(\tau)=1-\frac{2\lambda_{0}\,\alpha}{\mu\,(f(\theta_{0})-\tau\,\alpha)}\quad\text{;}\quad h^{\prime\prime}(\tau)=-\frac{2\lambda_{0}\,\alpha^{2}}{\mu\,(f(\theta_{0})-\tau\alpha)^{2}}$
Hence, $h(\tau)$ is maximized when at $\tau^{*}:=2{\lambda_{1}}{\mu}\,\left[f(\theta_{0})-\frac{\lambda_{0}}{\lambda_{1}}\right]$. Calculating $h(\tau^{*})$, we conclude that it is sufficient to set $T$ as:
$T\geq\frac{2}{\mu}\,\left[\lambda_{1}{f}(\theta_{0})+\lambda_{0}\,\left(\ln\left(\frac{\lambda_{0}}{\lambda_{1}\,\epsilon}\right)\right)\right]$
Appendix E Proofs for Section 6
See 3
Proof.
For an arbitrary comparator $u$ s.t. $f(u)\leq(\theta_{T})$, using the SGD update: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla{f_{t}}(\theta_{t})$,
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $=\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,\langle\nabla{f_{t}}(\theta_{t}),\theta_{t}-u\rangle+{\eta_{t}}^{2}\,\left\|\nabla{f_{t}}(\theta_{t})\right\|_{2}^{2}$ (26)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[{f_{t}}(\theta_{t})-{f_{t}}(u)]+{\eta_{t}}^{2}\,\left\|\nabla{f_{t}}(\theta_{t})\right\|_{2}^{2}$ (Convexity)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[{f_{t}}(\theta_{t})-{f_{t}}(u)]+\frac{{\eta_{t}}}{c}\,[{f_{t}}(\theta_{t})-{f_{t}}(\theta_{t+1})]$ (Using the stochastic Armijo line-search with $c>\frac{1}{2}$)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[{f_{t}}(\theta_{t})-{f_{t}}(u)]+\frac{{\eta_{t}}}{c}\,[{f_{t}}(\theta_{t})-{f_{t}}^{*}]$ (where ${f_{t}}^{*}:=\min{f_{t}}(\theta)$)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[{f_{t}}(\theta_{t})-{f_{t}}^{*}]-2{\eta_{t}}[{f_{t}}^{*}-{f_{t}}(u)]+\frac{{\eta_{t}}}{c}\,[{f_{t}}(\theta_{t})-{f_{t}}^{*}]$ (Add/subtract ${f_{t}}^{*}$)
Taking expectation w.r.t. the randomness at iteration $t$
$\implies\mathbb{E}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\mathbb{E}\left[{\eta_{t}}\underbrace{\left(2-\frac{1}{c}\right)}_{\text{Positive}}\,\underbrace{[{f_{t}}(\theta_{t})-{f_{t}}^{*}]}_{\text{Positive}}\right]+2\,\mathbb{E}\left[{\eta_{t}}\underbrace{[{f_{t}}(u)-{f_{t}}^{*}]}_{\text{Positive}}\right]$ (27)
From the line-search, we know that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{C^{\prime}}{{f_{t}}(\theta_{t})}\right\}$ where $C^{\prime}:=\frac{1}{\lambda_{1}}$.
Using this result, let us first upper-bound $-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$. For this we will consider two cases.
Case (1): If ${f_{t}}(\theta_{t})<f(\theta_{t})$,
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\min\left\{\eta_{\max}\,\mathbb{E}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\,,C^{\prime}\,\mathbb{E}\left[\frac{{f_{t}}(\theta_{t})-{f_{t}}^{*}}{{f_{t}}(\theta_{t})}\right]\right\}$ (Lower-bound on ${\eta_{t}}$)
$\leq-\min\left\{\eta_{\max}\,\mathbb{E}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right],C^{\prime}\,\mathbb{E}\left[\frac{{f_{t}}(\theta_{t})-{f_{t}}^{*}}{f(\theta_{t})}\right]\right\}$ (Using the case (1) condition)
$=-\min\left\{\eta_{\max},\frac{C^{\prime}}{f(\theta_{t})}\right\}\,\mathbb{E}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]$
$=-\underbrace{\min\left\{\eta_{\max},\frac{C^{\prime}}{f(\theta_{t})}\right\}}_{\text{Positive}}\,\bigg{[}\mathbb{E}\left[{f_{t}}(\theta_{t})-{f_{t}}(u)\right]+\underbrace{\mathbb{E}\left[{f_{t}}(u)-{f_{t}}^{*}\right]}_{\text{Positive}}\bigg{]}$ (Add/Subtract ${f_{t}}(u)$)
$\implies-\mathbb{E}\left[{\eta_{t}}\mathbb{E}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\min\left\{\eta_{\max},\frac{C^{\prime}}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]$
Case (2): If ${f_{t}}(\theta_{t})\geq f(\theta_{t})$,
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-{f_{t}}^{*}\right]\right]$ (Using the case (2) condition)
$=-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-{f_{t}}(u)\right]\right]-\underbrace{\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(u)-{f_{t}}^{*}\right]\right]}_{\text{Positive}}$ (Add/subtract ${f_{t}}(u)$)
$\implies-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]-\mathbb{E}\left[{\eta_{t}}\left[f(u)-{f_{t}}(u)\right]\right]$ (Add/subtract $f(u)$)
Let us consider two sub-cases: Case (i): If $f(u)-{f_{t}}(u)\geq 0$:
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]$
Case (ii): If $f(u)-{f_{t}}(u)<0$:
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]+\mathbb{E}\underbrace{\left[{\eta_{t}}\left[{f_{t}}(u)-f(u)\right]\right]}_{\text{Positive}}$
$\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]+\eta_{\max}\,\mathbb{E}\underbrace{\left[{f_{t}}(u)-f(u)\right]}_{=0}$
$\implies-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]$
Hence, in both sub-cases, we get that,
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\mathbb{E}\underbrace{\left[{\eta_{t}}\left[f(\theta_{t})-f(u)\right]\right]}_{\text{Positive}}\leq-\min\left\{\eta_{\max}\left[f(\theta_{t})-f(u)\right],C^{\prime}\,\mathbb{E}\left[\frac{f(\theta_{t})-f(u)}{{f_{t}}(\theta_{t})}\right]\right\}$ (Using the lower-bound on ${\eta_{t}}$)
$=-\min\left\{\eta_{\max}\left[f(\theta_{t})-f(u)\right],C^{\prime}\,[f(\theta_{t})-f(u)]\,\mathbb{E}\left[\frac{1}{{f_{t}}(\theta_{t})}\right]\right\}$ (Since $f(\theta_{t})-f(u)$ is independent of the randomness at iteration $t$)
$\leq-\min\left\{\eta_{\max}\left[f(\theta_{t})-f(u)\right],C^{\prime}\,[f(\theta_{t})-f(u)]\,\left[\frac{1}{\mathbb{E}[{f_{t}}(\theta_{t})]}\right]\right\}$ (Jensen’s inequality since $1/x$ is convex)
$\implies-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\min\left\{\eta_{\max}\left[f(\theta_{t})-f(u)\right],C^{\prime}\,\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]\right\}$ (Unbiasedness)
Hence, in both cases, we get that,
$-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\min\left\{\eta_{\max},\frac{C^{\prime}}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]$
Combining the above relations,
$\mathbb{E}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]+2\mathbb{E}\left[{\eta_{t}}\underbrace{[{f_{t}}(u)-{f_{t}}^{*}]}_{>0}\right]$ (where $C:=C^{\prime}\,\frac{2c-1}{c}$)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]+2\eta_{\max}\,\underbrace{\mathbb{E}[{f_{t}}(u)-{f_{t}}^{*}]}_{:=\chi^{2}(u)}$
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]+2\eta_{\max}\,\chi^{2}(u)$
Taking an expectation w.r.t the randomness iterations $t=0$ to $T-1$, and recursing, we conclude that
$\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]+2\eta_{\max}\,\chi^{2}(u)\,T$
See 5
Proof.
Define $u^{*}$ to be the max-margin solution i.e. $\left\|u^{*}\right\|=1$ and $\gamma$ to be the corresponding margin, i.e.
$\gamma$ $:=\min_{i}\,y_{i}\langle x_{i},u^{*}\rangle$ (28)
For a scalar $\beta>0$,
$f(\beta u^{*})$ $=\frac{1}{n}\,\sum_{i=1}^{n}\ln(1+\exp(-y_{i}\langle x_{i},\beta u^{*}\rangle))\leq\frac{1}{n}\,\sum_{i=1}^{n}\exp(-y_{i}\langle x_{i},u^{*}\rangle)\leq\exp(-\beta\gamma)$ (29)
For normalized data, s.t. $\left\|x_{i}\right\|=1$, the logistic regression loss is convex, is uniform smooth with $L=\lambda_{\max}[X^{T}X]=1$, satisfies Assumption 2 with $L_{1}=1$.
Consider $u=\beta u^{*}$ where $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon^{2}}\right)$ implies that $f(u)\leq\epsilon^{2}$. Since $\epsilon\leq 1$, $f(u)\leq\epsilon$.
For the $T$ defined in the theorem statement, consider two cases:
Case (I): $\mathbb{E}[f(\theta_{T})]<f(u)\leq\epsilon$. This gives the desired result immediately.
Case (II): $\mathbb{E}[f(\theta_{T})]>f(u)$. In this case, we can use the result in Lemma 3. In particular, for the comparator $u=\beta u^{*}$ where $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon^{2}}\right)$, SGD with stochastic Armijo line-search with $c$, and $\theta_{0}=0$ ensures that if $T$ is the first iteration s.t. $\mathbb{E}[f(\theta_{T})]-f(u)\leq\epsilon\implies\mathbb{E}[f(\theta_{T})]\leq\epsilon\,(1+\epsilon)$, then, for $C:=\frac{(2c-1)}{c\,\lambda_{1}}$, we conclude that,
$\mathbb{E}[f(\theta_{T})]-f(u)$ $\leq\frac{L}{2}\,\mathbb{E}\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{L}{2}\left[\left\|\theta_{0}-u\right\|_{2}^{2}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]+2\eta_{\max}\,\chi^{2}(u)\,T\right]$
$\implies\mathbb{E}[f(\theta_{T})]$ $\leq f(u)+\frac{L}{2}\left[\left\|u\right\|_{2}^{2}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]+2\eta_{\max}\,f(u)\,T\right]$ (Since $\chi^{2}(u)\leq f(u)$)
Setting $\eta_{\max}=\frac{C}{\epsilon}$ ensures that $\eta_{\max}\geq\frac{C}{f(\theta_{t})}$ for all $t\leq T$. Hence,
$\leq f(u)+\frac{L}{2}\left[\left\|u\right\|_{2}^{2}-C\,\mathbb{E}\sum_{t=0}^{T-1}\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]+\frac{2\,C\,f(u)\,T}{\epsilon}\right]$
Using that $f(u)\leq\epsilon^{2}$ and $\left\|u\right\|=\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon^{2}}\right)$,
$\implies\mathbb{E}[f(\theta_{T})]$ $\leq\epsilon^{2}+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}-C\,\mathbb{E}\sum_{t=0}^{T-1}\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]+2\,C\,\epsilon\,T\right]$
We also know that for all $t\leq T$, $f(\theta_{t})-f(u)\geq\epsilon$, meaning that $\mathbb{E}\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]\geq\frac{\epsilon}{f(u)+\epsilon}$. Using this relation,
$\leq\epsilon^{2}+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}-C\,T\,\left[\frac{\epsilon}{f(u)+\epsilon}\right]+2\,C\,\epsilon\,T\right]$
$\leq\epsilon^{2}+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}-\frac{C\,T}{2}+2\,C\,\epsilon\,T\right]$ (Since $f(u)\leq\epsilon^{2}\leq\epsilon$)
$\mathbb{E}[f(\theta_{T})]$ $\leq\epsilon^{2}+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}-\frac{3C\,T}{8}\right]$ (Since $\epsilon\leq\frac{1}{8}$)
Hence, in order to ensure that $f(\theta_{T})\leq\epsilon\,(1+\epsilon)\leq 2\epsilon$, it is sufficient to set $T$ to ensure that,
$\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}-\frac{3C\,T}{8}\right]\leq\epsilon$
Hence, it is sufficient to set $T$ as
$T$ $\geq\frac{8}{3C}\,\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}=\frac{16\,c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}$ (where we used $\lambda_{1}=\frac{6}{1-c}$ in $C$)